### v2

Abstract
Inexact proximal augmented Lagrangian methods (pALMs) are particularly appealing for tackling convex constrained optimization problems because of their elegant convergence properties and strong practical performance. To solve the associated pALM subproblems, efficient methods such as Newton-type methods are essential. Consequently, the effectiveness of the inexact pALM hinges on the error criteria used to control the inexactness when solving these subproblems. However, existing inexact pALMs either rely on absolute-type error criteria (which may complicate implementation by necessitating the pre-specification of an infinite sequence of error tolerance parameters) or require an additional correction step to guarantee convergence when using relative-type error criteria (which can potentially degrade the practical performance of pALM). To address these deficiencies, this paper proposes ripALM, a relative-type inexact pALM, for linearly constrained convex optimization. This method can simplify practical implementation while preserving the appealing convergence properties of the classical absolute-type inexact pALM. To the best of our knowledge, ripALM is the first relative-type inexact version of the vanilla pALM with provable convergence guarantees. Numerical experiments on quadratically regularized optimal transport problems and basis pursuit denoising problems demonstrate the competitive efficiency of the proposed method compared to existing state-of-the-art methods. As our model encompasses a wide range of application problems, the proposed ripALM offers broad applicability and has the potential to serve as a basic optimization toolbox.
Keywords: augmented Lagrangian method; proximal term; relative-type error criterion; asymptotic (super)linear convergence rate; quadratically regularized optimal transport; basis pursuit denoising
1 Introduction
Constrained optimization, admitting excellent modeling power for real-world applications across a wide range of fields, including machine learning and data science, engineering, operations research, is a central area in optimization, especially in the current big-data era [7, 10]. This paper is dedicated to solving the following linearly constrained convex optimization problem:
$\min_{\bm{x}\in\mathbb{R}^{N}}\quad f(\bm{x}),\quad\mathrm{s.t.}\quad A\bm{x}=\bm{b},$ (1.1)
where $f:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$ is a (possibly nonsmooth) proper closed convex function, $A\in\mathbb{R}^{M\times N}$ and $\bm{b}\in\mathbb{R}^{M}$ are given data. Problem (1.1) is a classical and widely studied template in convex optimization. Moreover, since the smoothness of the objective function $f$ is not required in this work, problem (1.1) actually encompasses a much broader class of application problems than it may appear at first glance. For instance, by leveraging an indicator function of a closed convex set, problem (1.1) naturally accommodates more general convex constraints; see also two illustrative examples in Section 4.
The augmented Lagrangian method (ALM) is recognized as one of the most popular and effective methods for solving constrained optimization problems [19, 37]. The essential component of the ALM for solving problem (1.1) involves penalizing the linear constraint $A\bm{x}=\bm{b}$ to derive the augmented Lagrangian function, defined as
$\mathcal{L}_{\sigma}^{\rm prim}(\bm{x},\bm{y})=f(\bm{x})+\left\langle\bm{y},A\bm{x}-\bm{b}\right\rangle+\frac{\sigma}{2}\left\lVert A\bm{x}-\bm{b}\right\rVert^{2},\quad(\bm{x},\bm{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M},$
where $\bm{y}\in\mathbb{R}^{M}$ is the Lagrangian multiplier associated with the linear constraint and $\sigma>0$ is a penalty parameter. Then, for a given sequence of penalty parameters $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$ and an initial Lagrangian multiplier $\bm{y}^{0}\in\mathbb{R}^{M}$, the ALM iteratively performs the following steps:
$\begin{cases}\bm{x}^{k+1}\in\arg\min\limits_{\bm{x}\in\mathbb{R}^{N}}\left\{\mathcal{L}_{\sigma_{k}}^{\rm prim}(\bm{x},\bm{y}^{k})\right\},\\[5.0pt] \bm{y}^{k+1}=\bm{y}^{k}+\sigma_{k}\nabla_{\bm{y}}\mathcal{L}_{\sigma_{k}}^{\rm prim}(\bm{x}^{k+1},\bm{y}^{k})=\bm{y}^{k}+\sigma_{k}\left(A\bm{x}^{k+1}-\bm{b}\right),\end{cases}$
which consists of minimizing one ALM-subproblem to get $\bm{x}^{k+1}$ and one gradient ascent step (associated with $\mathcal{L}_{\sigma_{k}}^{\rm prim}$) with a step size $\sigma_{k}$ to get $\bm{y}^{k+1}$. Since the augmented Lagrangian function $\mathcal{L}_{\sigma}^{\rm prim}$ is associated with the primal problem (1.1), the above ALM is usually named as a primal-based ALM. Alternatively, one may also consider a dual-based ALM which is associated with the dual problem of (1.1) given by (modulo a minus sign)
$\min\limits_{\bm{y}\in\mathbb{R}^{M}}\quad f^{*}({A}^{\top}\bm{y})-\bm{b}^{\top}\bm{y},$ (1.2)
where $f^{*}:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$ denotes the conjugate function of $f$ and $A^{\top}$ is the transpose of $A$. Then, given a penalty parameter $\sigma>0$, the augmented Lagrangian function associated with problem (1.2) is given by (see Section 2 for the detailed derivation)
$\mathcal{L}_{\sigma}^{\rm dual}(\bm{y},\bm{x})=-\bm{b}^{\top}\bm{y}+\frac{1}{2\sigma}\big\|\bm{x}+\sigma{A}^{\top}\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bm{x}\|^{2}-\mathtt{M}_{\sigma f}\big(\bm{x}+\sigma{A}^{\top}\bm{y}\big),~~(\bm{x},\bm{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M},$
where $\mathtt{M}_{\sigma f}(\bm{x}):=\min\limits_{\bm{z}}\left\{f(\bm{z})+\frac{1}{2\sigma}\|\bm{z}-\bm{x}\|^{2}\right\}$ denotes the Moreau envelope of the function $\sigma f$ at $\bm{x}$. Consequently, for a given sequence of penalty parameters $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$ and an initial point $\bm{x}^{0}\in\mathbb{R}^{N}$, the dual-based ALM reads as follows:
$\begin{cases}\bm{y}^{k+1}\in\arg\min\limits_{\bm{y}\in\mathbb{R}^{M}}\left\{\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y},\bm{x}^{k})\right\},\\[5.0pt] \bm{x}^{k+1}=\bm{x}^{k}+\sigma_{k}\nabla_{\bm{x}}\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y}^{k+1},\bm{x}^{k})=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big).\end{cases}$ (1.3)
The choice between using a primal-based or dual-based approach is problem-dependent. For example, when $f$ is nonsmooth, but its proximal mapping is easy-to-compute, the dual-based ALM may be more preferable. This is because, in the dual-based ALM, the optimization problem for updating $\bm{y}$ involves minimizing a continuously differentiable objective function, which can substantially simplify the optimization process.
The applicability of the ALM has been significantly expanded since Rockafellar’s seminal works [41, 42], which established a deep connection between the ALM and the proximal point algorithm (PPA). Moreover, the asymptotic (super)linear convergence rate of the ALM can be established under the assumption that the inverse of the subdifferential of the primal or dual objective function is Lipschitz continuous at the origin. Subsequent works have shown that this stringent condition can be further relaxed; see, for example, [11, 32] and references therein. Indeed, understanding the convergence properties of the ALM continues to be a topic of interest in the literature. Apart from theoretical advancements, numerous studies have demonstrated the remarkable efficiency of the ALM in solving various convex problems, including conic programming problems [23, 25, 27, 62], statistical optimization problems [22, 28, 60], optimization problems in machine learning [50, 54], and image/signal processing [21, 30], to mention just a few. These successful applications highlight the crucial roles of both the subsolvers employed to handle ALM subproblems and the error criteria used to control the accuracy required for solving these subproblems, while maintaining the desirable convergence properties of the ALM.
To solve ALM-subproblems efficiently, second-order methods, such as Newton-type methods [38], are strong candidates due to their fast local convergence rates. For example, once the iterate of a Newton-type method falls into the fast convergence region, only a few more iterations are needed to achieve a highly accurate solution. However, certain regularity conditions are needed to ensure the nonsingularity of the (generalized) Hessian of the ALM subproblem’s objective function at its solution point. Theoretically, such regularity conditions are generally not verifiable in advance, except in some special cases [27]. Numerically, ill-conditioned or singular (generalized) Hessians can lead to unpredictable numerical behaviors, resulting in unstable implementations and excessive efforts for parameter tuning. To enhance the applicability of a second-order method for solving ALM’s subproblems, a common simple approach is to add a proximal term to the ALM subproblem’s objective function, resulting in the proximal augmented Lagrangian method (pALM) [^1] The pALM is also known as the proximal method of multipliers; see [41].. Specifically, given $\{\tau_{k}\}\subseteq\mathbb{R}_{++}$, $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$ and $(\bm{x}^{0},\bm{y}^{0})\in\mathbb{R}^{N}\times\mathbb{R}^{M}$, the dual-based pALM iteratively performs the following steps:
$\begin{cases}\bm{y}^{k+1}=\arg\min\limits_{\bm{y}\in\mathbb{R}^{M}}\Big\{\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y},\bm{x}^{k})+\cfrac{\tau_{k}}{2\sigma_{k}}\left\|\bm{y}-\bm{y}^{k}\right\|^{2}\Big\},\\[8.0pt] \bm{x}^{k+1}=\bm{x}^{k}+\sigma_{k}\nabla_{\bm{x}}\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y}^{k+1},\bm{x}^{k})=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big).\end{cases}$ (1.4)
By incorporating a proximal term, the pALM is able to bypass the regularity conditions when applying a second-order method for solving its subproblem. This typically results in more robust implementations with reduced parameter tuning. We refer readers to [12, 13, 18, 20, 24, 26, 28, 29, 36, 58] for successful applications of the pALM in addressing various important optimization problems.
To make the pALM truly implementable and practical, it must allow approximate solutions to the subproblem with progressively improved accuracy, and the associated error criterion must be practically verifiable, while preserving desirable convergence properties of the outer loop. To achieve this, an absolute-type error criterion was introduced by Rockafellar for the pALM in [41] and is now widely used in the literature; see, for example, [24, 26, 28, 29, 36]. The term “absolute” means that the sequence of error tolerances, which controls the accuracy required for solving each pALM subproblem, is predetermined and remains fixed throughout the iterations. However, such a sequence typically needs to be summable and must be carefully tuned to avoid being either overly conservative or excessively aggressive. Consequently, achieving desirable numerical performance often demands delicate hyperparameter tuning, which may incur considerable effort and potential inefficiencies in practical implementations. A feasible alternative is to adopt a relative-type error criterion. The term “relative” refers to controlling the errors in minimizing the subproblems based on certain computable quantities related to the current status of the algorithm. This relative-type criterion eliminates the need for a summable tolerance parameter sequence, thereby enhancing the practical implementability of the algorithm. The adoption of a relative-type error criterion started from the seminal works of Solodov and Svaiter [45, 46, 47, 48, 49] on inexact versions of PPA, and has since influenced the development of inexact versions of numerous algorithms, including the variable metric PPA [33], ALM [14], ADMM [15], FISTA [4], and the Bregman-type methods [56, 59]. However, to the best of our knowledge, research on relative-type inexact versions of the pALM remains in its early stage, with limited exploration in [13, 20, 58]. In particular, all these inexact versions are derived by applying a (variable metric) hybrid proximal extragradient method [33, 45, 46] to a certain primal-dual solution mapping of the convex constrained problem, and therefore require an additional correction step to ensure convergence. Consequently, the algorithms developed in [13, 20, 58] essentially deviate from the vanilla pALM framework (1.4). Moreover, our numerical results also indicate that the correction step tends to slow down the convergence of the pALM and thus increase the computational cost. These observations naturally lead us to raise the following question:
Can we design a relative-type error criterion for the vanilla pALM, while preserving desirable convergence properties?
In light of this, the primary goal of this paper is to address the above question by designing a relative-type error criterion for the vanilla pALM framework (1.4). Note that the relative-type error criterion proposed by Eckstein and Silva [14] for the ALM innovatively introduces an auxiliary error variable to control the inexactness while ensuring convergence. In this work, we shall demonstrate that a similar idea can be adapted to the vanilla pALM, but with essential modifications. We should emphasize that the corresponding theoretical convergence analysis also requires substantial modifications, due to the inclusion of the proximal term. Interestingly, while the initial motivation for considering pALM was to ensure the nonsingularity of the (generalized) Hessian of the ALM subproblem’s objective function at its solution point, our analysis reveals that this proximal term can further improve the existing convergence properties of the relative-type inexact ALM studied in [1, 14, 61]. Specifically, under standard assumptions, it ensures the convergence of both primal and dual sequences, as well as their asymptotic (super)linear convergence rates. The key contributions and findings of this paper are summarized as follows:
  • We develop a relative-type inexact proximal augmented Lagrangian method (ripALM) to solve the dual problem (1.2). To the best of our knowledge, the proposed ripALM is the first inexact version of the vanilla pALM (1.4). It not only shares the same convergence properties for the sequence $\{\bm{x}^{k}\}$ as the relative-type inexact ALM studied in [1, 14, 61], but also offers additional theoretical advantages: the sequence $\{\bm{y}^{k}\}$ is guaranteed to be bounded; moreover, under a relatively weaker [^2] As noted in [24, Lemma 2.4], the error bound condition in Assumption A used in this work is relatively weaker than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin. The latter condition was used in [61] to establish the asymptotic (super)linear convergence rate for the relative-type inexact ALM. For this reason, we use the term “weaker” here. error bound assumption, the whole sequence $\{\bm{y}^{k}\}$ is further proved to be convergent, and both primal and dual sequences achieve asymptotic (super)linear convergence rates. These results provide stronger theoretical support for the applicability of ripALM, particularly in scenarios where accurate solutions of both primal and dual problems are required.
  • Numerically, we implement the proposed ripALM and apply it to solve quadratically regularized optimal transport (QROT) problems and basis pursuit denoising (BPDN) problems. Extensive numerical results validate the promising performance of the proposed ripALM for solving large-scale QROT and BPDN problems. Moreover, comparisons with existing methods underscore our motivation and contributions in developing a relative-type error criterion for the vanilla pALM. In addition, since problem (1.1) has strong modeling capability, our ripALM can potentially be extended to solve optimal transport problems with other convex regularizers such as the group quadratic regularizer [58], provided that the regularizers are proximal-friendly and the generalized Jacobians associated with the pALM-subproblems are easy to compute. Given the growing interests in developing efficient algorithms for computational optimal transport [34], our work may open up new possibilities for applying the powerful pALM algorithmic framework to a wide variety of real-world applications.
The remaining parts of this paper are organized as follows. Section 2 describes the main algorithmic framework of ripALM, whose convergence analysis is conducted in Section 3. Section 4 showcases how to apply the proposed ripALM for solving QROT and BPDN problems. Numerical experiments are conducted in Section 5. Finally, some concluding remarks are summarized in Section 6.
Notation. We use $\mathbb{R}^{n}$, $\mathbb{R}_{+}^{n}$, $\mathbb{R}^{m\times n}$ and $\mathbb{R}_{+}^{m\times n}$ to denote the sets of $n$-dimensional real vectors, $n$-dimensional non-negative vectors, $m\times n$ real matrices and $m\times n$ real non-negative matrices, respectively. For a vector $\bm{x}\in\mathbb{R}^{n}$, $x_{i}$ denotes its $i$-th entry, $\|\bm{x}\|$ denotes its Euclidean norm, and $\|\bm{x}\|_{H}:=\sqrt{\langle\bm{x},\,H\bm{x}\rangle}$ denotes its weighted norm associated with a symmetric positive definite matrix $H\in\mathbb{R}^{n\times n}$. For a matrix $X\in\mathbb{R}^{m\times n}$, $X_{ij}$} denotes its $(i,j)$-th entry, $\|X\|_{F}$ denotes its Frobenius norm, and $\texttt{vec}(X)$ denotes the vectorization of $X$, where $\texttt{vec}(X)]_{i+(j-1)m}=X_{ij}$ for any $1\leq i\leq m$ and $1\leq j\leq n$. For simplicity, given an integer $n>0$, we use $\bm{1}_{n}\in\mathbb{R}^{n}$ to denote the $n$-dimensional vector of all ones, and use $I_{n}$ to denote the $n\times n$ identity matrix.
For an extended-real-valued function $f:\mathbb{R}^{n}\rightarrow[-\infty,\infty]$, we say that it is proper if $f(\bm{x})>$-$\infty$ for all $\bm{x}\in\mathbb{R}^{n}$ and its effective domain ${\rm dom}\,f:=\{\bm{x}\in\mathbb{R}^{n}:f(\bm{x})<\infty\}$ is nonempty. A proper function $f$ is said to be closed if it is lower semicontinuous. Assume that $f:\mathbb{R}^{n}\rightarrow(-\infty,\infty]$ is a proper and closed convex function, the subdifferential of $f$ at $\bm{x}\in{\rm dom}\,f$ is defined by $\partial f(\bm{x}):=\big\{\bm{d}\in\mathbb{R}^{n}:f(\bm{y})\geq f(\bm{x})+\langle\bm{d},\,\bm{y}-\bm{x}\rangle,~\forall\,\bm{y}\in\mathbb{R}^{n}\big\}$ and its conjugate function $f^{*}:\mathbb{R}^{n}\rightarrow(-\infty,\infty]$ is defined by $f^{*}(\bm{y}):=\sup\big\{\langle\bm{y},\,\bm{x}\rangle-f(\bm{x}):\bm{x}\in\mathbb{R}^{n}\big\}$. For any $\nu>0$, the Moreau envelope of $\nu f$ at $\bm{x}$ is defined by $\mathtt{M}_{\nu f}(\bm{x}):=\min_{\bm{y}}\big\{f(\bm{y})+\frac{1}{2\nu}\|\bm{y}-\bm{x}\|^{2}\big\}$, and the proximal mapping of $\nu f$ at $\bm{x}$ is defined by $\mathtt{prox}_{\nu f}(\bm{x}):=\arg\min_{\bm{y}}\big\{f(\bm{y})+\frac{1}{2\nu}\|\bm{y}-\bm{x}\|^{2}\big\}$.
Let $\mathcal{S}$ be a closed convex subset of $\mathbb{R}^{n}$. Its indicator function $\delta_{\mathcal{S}}$ is defined by $\delta_{\mathcal{S}}(\bm{x})=0$ if $\bm{x}\in\mathcal{S}$ and $\delta_{\mathcal{S}}(\bm{x})=+\infty$ otherwise. Moreover, we denote the weighted distance of $\bm{x}\in\mathbb{R}^{n}$ to $\mathcal{S}$ by $\mathrm{dist}_{H}(\bm{x},\mathcal{S}):=\inf_{\bm{y}\in\mathcal{S}}\|\bm{x}-\bm{y}\|_{H}$ associated with a symmetric positive definite matrix $H$. When $H$ is the identity matrix, we omit $H$ in the notation and simply use $\mathrm{dist}(\bm{x},\mathcal{S})$ to denote the Euclidean distance of $\bm{x}\in\mathbb{R}^{n}$ to $\mathcal{S}$. Moreover, we use $\Pi_{\mathcal{S}}(\bm{x})$ to denote the projection of $\bm{x}$ onto $\mathcal{S}$.
2 A relative-type inexact pALM
In this section, we focus on developing a relative-type inexact proximal augmented Lagrangian method (ripALM) to solve the dual problem (1.2). The algorithmic framework is developed based on the parametric convex duality framework (see, for example, [39, 40] and [43, Chapter 11]).
We first identify problem (1.2) with the following problem
$\min\limits_{\bm{y}\in\mathbb{R}^{M}}\quad G(\bm{y},\bm{0}),$ (2.1)
where $G:\mathbb{R}^{M}\times\mathbb{R}^{N}\to(-\infty,+\infty]$ is defined by
$G(\bm{y},\bm{\xi}):=f^{*}({A}^{\top}\bm{y}+\bm{\xi})-\bm{b}^{\top}\bm{y}.$ (2.2)
Note that $G$ is proper closed convex since $f^{*}$ is proper closed convex. Then, the (ordinary) Lagrangian function of problem (1.2) can be defined by taking the concave conjugate of $G$ with respect to its second argument (see [43, Definition 11.45]), that is,
$\ell(\bm{y},\bm{x}):=\inf\limits_{\bm{\xi}\in\mathbb{R}^{N}}\left\{G(\bm{y},\bm{\xi})-\langle\bm{x},\,\bm{\xi}\rangle\right\}=-\bm{b}^{\top}\bm{y}+\langle\bm{x},\,{A}^{\top}\bm{y}\rangle-f(\bm{x}).$ (2.3)
Clearly, $\ell$ is convex in its first argument and concave in the second argument. For a given penalty parameter $\sigma>0$, the augmented Lagrangian function of problem (1.2) is defined by (see [43, Example 11.57])
$\mathcal{L}_{\sigma}(\bm{y},\,\bm{x})$= $\sup\limits_{\bm{s}\in\mathbb{R}^{N}}\left\{\ell(\bm{y},\bm{s})-\frac{1}{2\sigma}\|\bm{s}-\bm{x}\|^{2}\right\}$
$=$ $-\bm{b}^{\top}\bm{y}+\frac{1}{2\sigma}\big\|\bm{x}+\sigma{A}^{\top}\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bm{x}\|^{2}-\mathtt{M}_{\sigma f}\big(\bm{x}+\sigma{A}^{\top}\bm{y}\big).$
Here, we assume for the remainder of this paper that $\mathcal{L}_{\sigma}$ denotes the augmented Lagrangian function of the dual problem (1.2), omitting the superscript “dual” as in (1.3).
From the property of the Moreau envelope (see [2, Proposition 12.30]), we know that $\mathcal{L}_{\sigma}$ is continuously differentiable with respect to its first argument and
$\nabla_{y}\mathcal{L}_{\sigma}(\bm{y},\bm{x})=A\mathtt{prox}_{\sigma f}\big(\bm{x}+\sigma A^{\top}\bm{y}\big)-\bm{b}.$
With the above preparations, we are now ready to present our ripALM for solving problem (1.2) in Algorithm 1.
Algorithm 1 A relative-type inexact proximal augmented Lagrangian method (ripALM) for solving problem (1.2)
Input: $\rho\in[0,1)$, $\{\sigma_{k}\}_{k=0}^{\infty}$ such that $\inf_{k\geq 0}\{\sigma_{k}\}>0$, and $\{\tau_{k}\}_{k=0}^{\infty}$ such that $0<\inf_{k\geq 0}\{\tau_{k}\}\leq\sup_{k\geq 0}\{\tau_{k}\}<\infty$. Choose $\bm{y}^{0},\,\bm{w}^{0}\in\mathbb{R}^{M}$ and $\bm{x}^{0}\in\mathbb{R}^{N}$ arbitrarily. Set $k=0$.
while the termination criterion is not met, do
  • Step 1.
    Approximately solve the subproblem:
    $\min\limits_{\bm{y}\in\mathbb{R}^{M}}~~\mathcal{L}_{\sigma_{k}}(\bm{y},\,\bm{x}^{k})+\frac{\tau_{k}}{2\sigma_{k}}\big\|\bm{y}-\bm{y}^{k}\big\|^{2}$ (2.4)
    to find $\bm{y}^{k+1}\in\mathbb{R}^{M}$ such that
     $2\big|\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle\big|+\big\|\sigma_{k}\Delta^{k+1}\big\|^{2}$ (2.5)
    $\leq$ $\;\rho\left(\big\|\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big)-\bm{x}^{k}\big\|^{2}+\tau_{k}\big\|\bm{y}^{k+1}-\bm{y}^{k}\big\|^{2}\right),$
    where $\Delta^{k+1}:=\nabla_{y}\mathcal{L}_{\sigma_{k}}(\bm{y}^{k+1},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k+1}-\bm{y}^{k})$.
  • Step 2.
    Update
    $\bm{x}^{k+1}=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big),\quad\bm{w}^{k+1}=\bm{w}^{k}-\sigma_{k}\Delta^{k+1}.$ (2.6)
  • Step 3.
    Set $k=k+1$ and go to Step 1.
end while
Output: $(\bm{y}^{k},\bm{x}^{k})$.
In line with pALM-type methods, at each iteration, our ripALM in Algorithm 1 approximately minimizes the sum of the augmented Lagrangian function $\mathcal{L}_{\sigma_{k}}(\cdot,\,\bm{x}^{k})$ and a proximal term $\frac{\tau_{k}}{2\sigma_{k}}\big\|\cdot-\bm{y}^{k}\big\|^{2}$ under the error criterion (2.5), followed by the updates of the multiplier $\bm{x}^{k}$ and the auxiliary error variable $\bm{w}^{k}$. Next, we show that the error criterion (2.5) is achievable at each iteration, provided that a convergent subsolver is employed. Specifically, the subproblem (2.4) admits a unique solution $\bm{y}^{k,*}$ due to the strong convexity, which satisfies
$\nabla_{y}\mathcal{L}_{\sigma_{k}}(\bm{y}^{k,*},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k,*}-\bm{y}^{k})=\bm{0}.$
Then, a suitable subsolver (e.g., a gradient-type or Newton-type method) for solving the subproblem (2.4) can generate a sequence $\{\bm{y}^{k,t}\}$ converging to $\bm{y}^{k,*}$ as $t\to\infty$. Since the objective function in (2.4) is continuously differentiable (by the property of the Moreau envelope [43]), it follows that
$\Delta^{k,t}:=\nabla_{y}\mathcal{L}_{\sigma_{k}}(\bm{y}^{k,t},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k,t}-\bm{y}^{k})\to\bm{0}\quad\mathrm{as}\quad t\to\infty,$
which further implies that
$2\big|\langle\bm{w}^{k}-\bm{y}^{k,t},\,\sigma_{k}\Delta^{k,t}\rangle\big|+\big\|\sigma_{k}\Delta^{k,t}\big\|^{2}\to 0\quad\mathrm{as}\quad t\to\infty.$
Therefore, the left-hand side of (2.5) converges to 0 as $t\to\infty$. On the other hand, if $\bm{y}^{k}$ is not the unique solution of the subproblem (2.4) with $\tau_{k}\neq 0$ (which is the case of interest, as otherwise we can directly proceed to the next iteration), then the term $\tau_{k}\|\bm{y}^{k,t}-\bm{y}^{k}\|^{2}$ and thus the right-hand side of (2.5) would not approach 0 as $t\to\infty$. Consequently, in this case, (2.5) must be satisfied at some iterate $\bm{y}^{k,t}$ after finitely many $t$ inner iterations. Once (2.5) is satisfied at $\bm{y}^{k,t}$, we set $\bm{y}^{k+1}:=\bm{y}^{k,t}$ and proceed to the next outer iteration.
Compared to the recent semismooth Newton based inexact proximal augmented Lagrangian (Snipal) method developed in [24, Section 3], our ripALM in Algorithm 1 employs a significantly different error criterion (2.5) for solving the subproblem (2.4). Specifically, in our context, Snipal requires the error term $\Delta^{k+1}$ to satisfy
$(A)~~\big\|\Delta^{k+1}\big\|\leq\frac{\min\left(\sqrt{\tau_{k}},\,1\right)}{\sigma_{k}}\varepsilon_{k},~~\varepsilon_{k}\geq 0,~~\sum_{k=0}^{\infty}\varepsilon_{k}<\infty,$ (2.7)
$(B)~~\big\|\Delta^{k+1}\big\|\leq\frac{\delta_{k}\min\left(\sqrt{\tau_{k}},\,1\right)}{\sigma_{k}}\sqrt{\big\|\Delta_{x}^{k+1}\big\|^{2}+\tau_{k}\big\|\Delta_{y}^{k+1}\big\|^{2}},~~0\leq\delta_{k}<1,~~\sum_{k=0}^{\infty}\delta_{k}<\infty,$
with $\Delta_{x}^{k+1}:=\bm{x}^{k+1}-\bm{x}^{k}$ and $\Delta_{y}^{k+1}:=\bm{y}^{k+1}-\bm{y}^{k}$, to guarantee an asymptotic (super)linear convergence rate. [^3] Note that the error criterion (A) in (2.7) alone is sufficient for establishing the global convergence of the Snipal; see [24, Section 3]. Both error criteria (A) and (B) in (2.7) are of the absolute type, meaning that they require the pre-specification of two summable tolerance parameter sequences, $\{\varepsilon_{k}\}\subseteq[0,\infty)$ and $\{\delta_{k}\}\subseteq[0,1)$, to control the error incurred in the inexact minimization of the subproblem. Since there is generally no direct guidance on optimally selecting these tolerance parameters to achieve good convergence efficiency, this absolute-type criterion typically requires careful tuning for the tolerance parameters, which may make Snipal less user-friendly in practical implementations. In contrast, the error criterion (2.5) used in ripALM is of the relative type, meaning that the error $2\big|\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle\big|+\big\|\sigma_{k}\Delta^{k+1}\big\|^{2}$ is regulated by a tentative successive difference related to the progress of the algorithm. Moreover, this relative-type criterion only involves a single tolerance parameter $\rho\in[0,1)$, thus simplifying the process of parameter tuning in both computation and implementation, as we shall see in subsection 5.1.
Thanks to the advantage of eliminating the need to select an infinite sequence of tolerance parameters, relative-type error criteria have been widely adopted in numerous well-known algorithms (e.g., PPA, ALM, ADMM, FISTA, etc.) to approximately solve subproblems over the past two decades. This trend began with the seminal works of Solodov and Svaiter [45, 46, 47, 48, 49], which has since inspired the development of numerous relative-type inexact algorithms; see, for example, [4, 14, 15, 33, 56, 58, 59]. Following this line of research, Yang et al. [58] recently developed a corrected inexact proximal augmented Lagrangian method (ciPALM) for solving a class of group-quadratic regularized optimal transport problems. The error criterion used there can be described as follows: choose $\rho\in[0,1)$, at each iteration, approximately solve the subproblem (2.4) to find a point $\widetilde{\bm{y}}^{k+1}$ such that
$\widetilde{\Delta}^{k+1}$= $~\nabla_{y}\mathcal{L}_{\sigma_{k}}(\widetilde{\bm{y}}^{k+1},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\widetilde{\bm{y}}^{k+1}-\bm{y}^{k}),$ (2.8)
$\hskip-7.11317pt\big\|{\sigma_{k}}\widetilde{\Delta}^{k+1}\big\|^{2}\leq$ $~\rho^{2}\min\left({\tau_{k}},\,1\right)\left(\left\|\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\widetilde{\bm{y}}^{k+1}\big)-\bm{x}^{k}\right\|^{2}+\tau_{k}\left\|\widetilde{\bm{y}}^{k+1}-\bm{y}^{k}\right\|^{2}\right).$
When (2.8) is satisfied, the multiplier is updated as usual: $\bm{x}^{k+1}=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\widetilde{\bm{y}}^{k+1}\big)$. This is then followed by an extra correction step: $\bm{y}^{k+1}=\bm{y}^{k}-\tau_{k}^{-1}\sigma_{k}\left(A\bm{x}^{k+1}-\bm{b}\right)$. After these updates, the algorithm proceeds to the next iteration. One can see that the error criterion (2.8) used in ciPALM is also of the relative type and looks even simpler than (2.5) used in our ripALM. However, we should point out that ciPALM requires an additional correction step to ensure its convergence, and hence it may only be viewed as a pALM-like algorithm. Moreover, our numerical results indicate that this correction step tends to degrade the practical performance of the pALM framework. This raises the natural question of whether one can design a relative-type error criterion directly for the vanilla pALM framework. Addressing this question serves as the main motivation of this work. To this end, we propose the relative error criterion (2.5), where the error variable $\bm{w}^{k}$ is used solely in the construction of the error criterion (2.5) and does not directly influence either the objective function in (2.4) or the updating rules of the primal and dual variables. Our numerical results show that the proposed ripALM may offer greater robustness and efficiency; see subsection 5.1 for numerical comparisons.
Our relative error criterion (2.5) is inspired by Eckstein and Silva’s practical relative error criterion [14], which was developed for the approximate minimization of the subproblems in the vanilla ALM (i.e., without the proximal term $\frac{\tau_{k}}{2\sigma_{k}}\big\|\bm{y}-\bm{y}^{k}\big\|^{2}$ in the subproblem (2.4) in our context). Unlike their work, we suggest incorporating $\frac{\tau_{k}}{2\sigma_{k}}\big\|\bm{y}-\bm{y}^{k}\big\|^{2}$ in the subproblem (2.4). This proximal term guarantees the existence and uniqueness of the optimal solution of the strongly convex subproblem (2.4). Introducing the proximal term also ensures the positive definiteness of the generalized Hessian/Jacobian of the objective function in (2.4), thereby facilitating the application of the semi-smooth Newton method to the subproblem (2.4) effectively, as shown in Section 4.1. In addition, as we will see later in the next section, our ripALM not only shares the same convergence properties for the sequence $\{\bm{x}^{k}\}$ as Eckstein and Silva’s relative-type inexact ALM (denoted by ES-iALM) developed in [1, 14, 61], but also offers additional theoretical advantages: the sequence $\{\bm{y}^{k}\}$ is guaranteed to be bounded; moreover, under a relatively weaker error bound assumption (than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin), the sequence $\{\bm{y}^{k}\}$ is further proved to be convergent, and both primal and dual sequences achieve asymptotic (super)linear convergence rates.
3 Convergence analysis
In this section, we study the convergence properties of the ripALM in Algorithm 1. Before proceeding, we recall the definition of $G$ from (2.2) and further define $F:\mathbb{R}^{M}\times\mathbb{R}^{N}\rightrightarrows\mathbb{R}^{M}\times\mathbb{R}^{N}$ to be the concave conjugate of $G$:
$F(\bm{\theta},\bm{x}):=\inf\limits_{\bm{y}\in\mathbb{R}^{M},\,\bm{\xi}\in\mathbb{R}^{N}}\big\{G(\bm{y},\bm{\xi})-\langle\bm{\theta},\bm{y}\rangle-\langle\bm{x},\bm{\xi}\rangle\big\},$ (3.1)
which is a closed (upper semicontinuous) concave function. Then, the dual problem of (2.1) is given by
$\max\limits_{\bm{x}\in\mathbb{R}^{N}}\quad F(\bm{0},\bm{x}),$ (3.2)
which can be rewritten as problem (1.1). Next, let $\partial G:\mathbb{R}^{M}\times\mathbb{R}^{N}\rightrightarrows\mathbb{R}^{M}\times\mathbb{R}^{N}$ and $\partial F:\mathbb{R}^{M}\times\mathbb{R}^{N}\rightrightarrows\mathbb{R}^{M}\times\mathbb{R}^{N}$ denote the subgradient maps of $G$ and $F$, respectively, that is,
$(\bm{\theta},\bm{x})\in\partial G(\bm{y},\bm{\xi})~\Leftrightarrow~G(\bm{y}^{\prime},\bm{\xi}^{\prime})\geq G(\bm{y},\bm{\xi})+\langle\bm{\theta},\bm{y}^{\prime}-\bm{y}\rangle+\langle\bm{x},\bm{\xi}^{\prime}-\bm{\xi}\rangle,~~\forall(\bm{y}^{\prime},\bm{\xi}^{\prime}),$
$(\bm{y},\bm{\xi})\in\partial F(\bm{\theta},\bm{x})~\Leftrightarrow~F(\bm{\theta}^{\prime},\bm{x}^{\prime})\leq F(\bm{\theta},\bm{x})-\langle\bm{y},\bm{\theta}^{\prime}-\bm{\theta}\rangle-\langle\bm{\xi},\bm{x}^{\prime}-\bm{x}\rangle,~~\forall(\bm{\theta}^{\prime},\bm{x}^{\prime}).$
We also recall the definition of $\ell$ from (2.3) and let $\partial\ell:\mathbb{R}^{M}\times\mathbb{R}^{N}\rightrightarrows\mathbb{R}^{M}\times\mathbb{R}^{N}$ denote the subgradient map of $\ell$, that is,
$(\bm{\theta},\bm{\xi})\in\partial\ell(\bm{y},\bm{x})~\Leftrightarrow~\left\{\begin{aligned} &\ell(\bm{y}^{\prime},\bm{x})\geq\ell(\bm{y},\bm{x})+\langle\bm{\theta},\bm{y}^{\prime}-\bm{y}\rangle,~~\forall\bm{y}^{\prime}\in\mathbb{R}^{M},\\ &\ell(\bm{y},\bm{x}^{\prime})\leq\ell(\bm{y},\bm{x})-\langle\bm{\xi},\bm{x}^{\prime}-\bm{x}\rangle,~~\forall\bm{x}^{\prime}\in\mathbb{R}^{N}.\end{aligned}\right$.
By this definition, one can verify that
$\partial\ell(\bm{y},\bm{x})=\left\{A\bm{x}-\bm{b}\right\}\times\left\{\bm{v}-A^{\top}\bm{y}\mid\bm{v}\in\partial f(\bm{x})\right\}.$
Moreover, all the subgradient maps $\partial G$, $\partial F$, and $\partial\ell$ are maximal monotone operators, and satisfy that (see also [14, equation (23)]):
$(\bm{\theta},\bm{x})\in\partial G(\bm{y},\bm{\xi})~\Leftrightarrow~(\bm{\theta},\bm{\xi})\in\partial\ell(\bm{y},\bm{x})~\Leftrightarrow~(\bm{y},\bm{\xi})\in\partial F(\bm{\theta},\bm{x}).$ (3.3)
If $(\bm{y}^{*},\bm{x}^{*})\in\mathbb{R}^{M}\times\mathbb{R}^{N}$ satisfies $(\bm{0},\bm{0})\in\partial\ell(\bm{y}^{*},\bm{x}^{*})$, i.e, $(\bm{y}^{*},\bm{x}^{*})$ solves the following KKT system:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bm{0}\in\partial f(\bm{x})-A^{\top}\bm{y},\quad A\bm{x}-\bm{b}=\bm{0},$
then $\bm{y}^{*}$ solves problem (2.1) (i.e., problem (1.2)) and $\bm{x}^{*}$ solves problem (3.2) (i.e., problem (1.1)). In this case, we call $(\bm{y}^{*},\bm{x}^{*})$ a saddle point of the Lagrangian function $\ell(\bm{y},\bm{x})$. If such a saddle point exists, then the strong duality holds, that is, the optimal values of problems (2.1) and (3.2) are finite and equal, i.e., $G(\bm{y}^{*},\bm{0})=F(\bm{0},\bm{x}^{*})$. Moreover, the set of saddle points can be written as $\mathcal{Y}^{*}\times\mathcal{X}^{*}\subseteq\mathbb{R}^{M}\times\mathbb{R}^{N}$, where $\mathcal{Y}^{*}$ is the solution set of problem (2.1) (i.e., problem (1.2)) and $\mathcal{X}^{*}$ is the solution set of problem (3.2) (i.e., problem (1.1)).
With the above preparations, we are now ready to establish the convergence results of the proposed ripALM in Algorithm 1. Note that, while our analysis is inspired by that of ES-iALM in [1, 14, 61], the inclusion of the proximal term $\frac{\tau_{k}}{2\sigma_{k}}\|\bm{y}-\bm{y}^{k}\|^{2}$ indeed introduces new technical challenges. We need to carefully derive new recursive inequalities that simultaneously involve both primal and dual sequences, $\{\bm{x}^{k}\}$ and $\{\bm{y}^{k}\}$, which cannot be obtained by applying existing results directly. In particular, to establish the asymptotic (super)linear convergence rate in Theorem 3.2, we derive a novel recursion (A.20), which critically relies on the presence of a strictly positive proximal parameter $\tau_{k}$, as $\tau_{k}$ appears in the denominator of the coefficient on the right-hand side. Consequently, this novel recursion is tailored to our ripALM framework.
Theorem 3.1.
Let the functions $G$, $F$ and $\ell$ be defined as in (2.2), (3.1) and (2.3), respectively. Let $\rho\in[0,1)$, $\{\sigma_{k}\}$ be a positive sequence satisfying that $\sigma_{k}\geq\sigma_{\min}>0$ for all $k\geq 0$, and $\{\tau_{k}\}$ be a positive sequence satisfying that
$\tau_{k}\geq\tau_{\min}>0,\quad\tau_{k+1}\leq(1+\nu_{k})\tau_{k}\quad\mbox{with}\quad\nu_{k}\geq 0~~\mbox{and}~~{\textstyle\sum_{k=0}^{\infty}}\nu_{k}<+\infty.$
Let $\{\bm{y}^{k}\}$, $\{\Delta^{k}\}$, $\{\bm{w}^{k}\}\subset\mathbb{R}^{M}$ and $\{\bm{x}^{k}\}\subset\mathbb{R}^{N}$ be sequences generated by Algorithm 1. If $\ell$ admits a saddle point (i.e., $(\partial\ell)^{-1}(\bm{0},\bm{0})\neq\emptyset$), then the following statements hold.
  1. (i)
    The sequences $\{\bm{y}^{k}\}$, $\{\bm{w}^{k}\}$ and $\{\bm{x}^{k}\}$ are bounded.
  2. (ii)
    $\lim\limits_{{k}\to\infty}\Delta^{k+1}=\bm{0}$, $\lim\limits_{{k}\to\infty}\bm{\theta}^{k+1}=\bm{0}$ and $\lim\limits_{{k}\to\infty}\bm{\xi}^{k+1}=\bm{0}$, where $\bm{\theta}^{k+1}$ and $\bm{\xi}^{k+1}$ are defined by
    $\bm{\theta}^{k+1}:=\Delta^{k+1}-\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k+1}-\bm{y}^{k})\quad\text{and}\quad\bm{\xi}^{k+1}:=\sigma_{k}^{-1}(\bm{x}^{k}-\bm{x}^{k+1}),\quad\forall\,k\geq 0.$
  3. (iii)
    Both sequences $\left\{G(\bm{y}^{k+1},\bm{\xi}^{k+1})\right\}$ and $\left\{F(\bm{\theta}^{k+1},\,\bm{x}^{k+1})\right\}$ converge to the common optimal value of problems (2.1) and (3.2).
  4. (iv)
    Any accumulation point of $\{\bm{y}^{k}\}$ is an optimal solution of problem (2.1) (i.e., problem (1.2)), and any accumulation point of $\{\bm{x}^{k}\}$ is an optimal solution of problem (3.2) (i.e., problem (1.1)).
  5. (v)
    The sequence $\{\bm{x}^{k}\}$ converges to an optimal solution of problem (3.2).
Proof.
See Appendix A.1.    $\Box$
Next, we study the asymptotic (super)linear convergence rate of our ripALM under an error bound condition presented in Assumption A. As noted in [24, Lemma 2.4], this error-bound condition is weaker than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin. The latter condition was used in [61] to establish the asymptotic (super)linear convergence rate for Eckstein and Silva’s relative-type inexact ALM, while the former, weaker condition has also been employed in [24] and [58] to establish the asymptotic (super)linear convergence rate for Snipal and ciPALM, respectively.
Assumption A.
For any $r>0$, there exists a constant $\kappa>0$ such that, for any $(\bm{y},\bm{x})\in\left\{(\bm{y},\bm{x})\in\mathbb{R}^{M}\times\mathbb{R}^{N}\mid\mathrm{dist}\left((\bm{y},\bm{x}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right)\leq r\right\}$,
$\mathrm{dist}\left((\bm{y},\bm{x}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right)\leq\kappa\,\mathrm{dist}\left((\bm{0},\bm{0}),\,\partial\ell(\bm{y},\bm{x})\right).$ (3.4)
Theorem 3.2.
Let $\ell$ be defined as in (2.3), and let $\rho\in[0,1)$, $\{\sigma_{k}\}$ be a positive sequence satisfying that $\sigma_{k}\geq\sigma_{\min}>0$ for all $k\geq 0$, and $\{\tau_{k}\}$ be a positive sequence satisfying that
$\tau_{k}\geq\tau_{\min}>0,\quad\tau_{k+1}\leq(1+\nu_{k})\tau_{k}\quad\mbox{with}\quad\nu_{k}\geq 0~~\mbox{and}~~{\textstyle\sum_{k=0}^{\infty}}\nu_{k}<+\infty.$
Suppose additionally that $\ell$ admits a saddle point (i.e., $(\partial\ell)^{-1}(\bm{0},\bm{0})\neq\emptyset$), Assumption A holds, and the sequences of parameters $\rho$, $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ satisfy that
$\sqrt{\tau_{\min}}-2\sqrt{\rho}>0\quad\text{and}\quad\liminf\limits_{k\to\infty}~\sigma_{k}>c\cdot\frac{2\kappa\sqrt{\tau_{\max}}\left(\rho+\sqrt{\rho\,\overline{\tau}_{\max}}\right)}{\sqrt{\tau_{\min}}-2\sqrt{\rho}},$ (3.5)
where $c>1$ is an arbitrarily given positive constant, $\tau_{\max}:=\tau_{0}\prod_{k=0}^{\infty}(1+\nu_{k})$, and $\overline{\tau}_{\max}:=\max\left\{1,\tau_{\max}\right\}$. Let $\Lambda^{k}:=\operatorname{Diag}(\tau_{k}I_{M},I_{N})$, $\overline{\tau}_{k}:=\max\left\{1,\tau_{k}\right\}$ and
$\gamma_{k}:=\left(1-\frac{2\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)+2\sigma_{k}\sqrt{\rho}}{\sigma_{k}\sqrt{\tau_{k}}}\right)\frac{\sigma_{k}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{k}}\right)^{2}\overline{\tau}_{k}}.$
Then, the following statements hold.
  1. (i)
    For all sufficiently large $k$, we have that
    $\gamma_{k}\geq\left(\frac{c-1}{c}\right)\cdot\frac{\sqrt{\tau_{\min}}-2\sqrt{\rho}}{\sqrt{\tau_{\min}}}\cdot\frac{\sigma_{k}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{\max}}\right)^{2}\overline{\tau}_{\max}}>0,$
    and
    $\mathrm{dist}_{\Lambda^{k+1}}\left((\bm{y}^{k+1},\bm{x}^{k+1}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right)\leq\mu_{k}\,\mathrm{dist}_{\Lambda^{k}}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right),$
    where
    $\mu_{k}:=\sqrt{\frac{1+\nu_{k}}{1+\gamma_{k}}}\quad\mbox{satisfies}\quad\limsup\limits_{k\to\infty}\,\left\{\mu_{k}\right\}<1~~\mbox{as}~~\nu_{k}\to 0.$
  2. (ii)
    The sequence $\{\bm{y}^{k}\}$ is convergent.
Proof.
See Appendix A.2.    $\Box$
Remark 3.1 (Comments on $\gamma_{k}$ and $\mu_{k}$).
One can see from the expression of $\gamma_{k}$ in Theorem 3.2 that, after a finite number of iterations, $\gamma_{k}$ becomes proportional to the squared penalty parameter $\sigma_{k}^{2}$, provided that $\rho$, $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ satisfy the conditions in (3.5). Therefore, if $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$, the convergence factor $\mu_{k}$ can be less than 1 when $\sigma_{k}$ is sufficiently large, and it approaches 0 as $\sigma_{k}$ tends to $\infty$. This readily demonstrates that the sequence $\left\{(\bm{y}^{k},\bm{x}^{k})\right\}$ converges to the set of saddle points at a Q-(super)linear rate if $\sigma_{k}$ is sufficiently large and $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$. In practical implementations, one could simply choose an increasing sequence of $\{\sigma_{k}\}$ with $\sigma_{k}\uparrow+\infty$, and set $\tau_{\min}\geq 4$ to ensure that $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$ for any $\rho\in[0,1)$. In contrast, as discussed in [58, Remark 2.1], ciPALM (which is also of relative-type but uses the error criterion (2.8)) should require $\rho<\frac{1}{3}$ to guarantee an asymptotic (super)linear convergence rate under the same error bound condition in Assumption A. This highlights another advantage of our ripALM, as it offers greater flexibility in choosing the tolerance parameter $\rho$.
Before closing this section, we summarize in Table 1 a brief comparison among the related inexact (p)ALMs, based on the discussions in Section 2 and the convergence results established in this section. From the comparison, we see that ripALM enjoys stronger theoretical guarantees than ES-iALM and, unlike ciPALM, does not require correction steps.
Table 1: A brief comparison among ES-iALM, Snipal, ciPALM, and ripALM. The convergence results for each algorithm are presented in the context of solving problem (1.2), which is consistent with our setting. In the table, “Seq. Conv.”: sequential convergence; “Abs.”: absolute-type; “Rel.”: relative-type; ✔E: be convergent under error bound condition.
Methods Type
Vanilla
(p)ALM
Proximal
Term
Boundedness Seq. Conv.
Asymptotic
(super)linear Rate
(under Error Bound)
$\{\bm{y}^{k}\}$ $\{\bm{x}^{k}\}$ $\{\bm{y}^{k}\}$ $\{\bm{x}^{k}\}$
ES-iALM [1, 14, 61] Rel. $\{\bm{x}^{k}\}$
Snipal [24] Abs. $\{(\bm{y}^{k},\bm{x}^{k})\}$
ciPALM [58] Rel.
$\{(\bm{y}^{k},\bm{x}^{k})\}$
ripALM [this work] Rel. E $\{(\bm{y}^{k},\bm{x}^{k})\}$
4 Applications of ripALM
In this section, we present the application of the ripALM in Algorithm 1 to two important problems: the quadratically regularized optimal transport problem and the basis pursuit denoising problem.
4.1 Quadratically regularized optimal transport
As an important variant of the classical optimal transport problem, the quadratically regularized optimal transport (QROT) problem introduces a quadratic regularization (i.e., ridge regularization) term into the objective function. This renders the problem strongly convex while preserving the sparsity of the optimal transport plan. These favorable properties make QROT a compelling alternative to the popular entropically regularized optimal transport problem, especially in applications where the sparse structure is desirable; see, e.g., [5, 16, 31] for more details. Mathematically, the QROT problem is given as follows:
$\min\limits_{X}~~\frac{\lambda}{2}\|X\|_{F}^{2}+\langle C,\,X\rangle\quad\mathrm{s.t.}\quad X\bm{1}_{n}=\bm{\alpha},~~X^{\top}\bm{1}_{m}=\bm{\beta},~~X\geq 0,$ (4.1)
where $\lambda>0$ is a given regularization parameter, $C\in\mathbb{R}^{m\times n}_{+}$ is a given cost matrix, $\bm{\alpha}:=(\alpha_{1},\cdots,\alpha_{m})^{\top}\in\Sigma_{m}$ and $\bm{\beta}:=(\beta_{1},\cdots,\beta_{n})^{\top}\in\Sigma_{n}$ are given probability vectors with $\Sigma_{m}$ (resp., $\Sigma_{n}$) denoting the $m$ (resp., $n$)-dimensional unit simplex, and $\bm{1}_{m}$ (resp., $\bm{1}_{n}$) denotes the $m$ (resp., $n$)-dimensional vector of all ones.
To apply the proposed ripALM, we first reformulate (4.1) as follows:
$\min_{X\in\mathbb{R}^{m\times n}}~~f_{\texttt{q}}(X):=\frac{\lambda}{2}\left\|{X}\right\|_{F}^{2}+\left\langle{C},{X}\right\rangle+\delta_{\mathbb{R}^{m\times n}_{+}}(X)\quad\mathrm{s.t.}\quad X\bm{1}_{n}=\bm{\alpha},~X^{\top}\bm{1}_{m}=\bm{\beta}.$ (4.2)
Clearly, it falls into the form of problem (1.1) (upon the vectorization of $X$). Further, one can show that the dual problem of (4.2) is given by (modulo a minus sign)
$\min\limits_{\bm{u},\,\bm{v}}~~f_{\texttt{q}}^{*}\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v},$ (4.3)
and its associated augmented Lagrangian function is given by (using a similar deduction as Section 2)
$\mathcal{L}_{\sigma}(\bm{u},\bm{v},X)=$ $-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v}+\frac{1}{2\sigma}\left\|X+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right\|_{F}^{2}$
$-\frac{1}{2\sigma}\|X\|_{F}^{2}-\mathtt{M}_{\sigma f_{\texttt{q}}}\left(X+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right),$
where $\bm{u}\in\mathbb{R}^{m}$, $\bm{v}\in\mathbb{R}^{n}$ are the Lagrange multipliers corresponding to $X\bm{1}_{n}=\bm{\alpha}$ and $X^{\top}\bm{1}_{m}=\bm{\beta}$, respectively, and the conjugate function $f_{\texttt{q}}^{*}$ admits the following expression:
$f_{\texttt{q}}^{*}(Z)=\left\{\begin{aligned} &\delta_{\mathbb{R}^{m\times n}_{+}}\left(C-Z\right),&&\mathrm{if}~~\lambda=0,\\[3.0pt] &\frac{1}{2\lambda}\left\|\Pi_{\mathbb{R}^{m\times n}_{+}}\left(Z-C\right)\right\|_{F}^{2},&&\mathrm{if}~~\lambda>0.\end{aligned}\right$.
Thus, given an arbitrary initial guess $\left(\bm{u}^{0},\bm{v}^{0},X^{0}\right)\in\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{m\times n}$, the basic iterative scheme of our ripALM for solving (4.3) reads as follows:
$\left(\bm{u}^{k+1},\bm{v}^{k+1}\right)\approx\mathop{\arg\min}\limits_{\bm{u}\in\mathbb{R}^{m},\bm{v}\in\mathbb{R}^{n}}\left\{\Psi_{k}(\bm{u},\bm{v})\right\},$ (4.4)
$X^{k+1}=\mathtt{prox}_{\sigma_{k}f_{\texttt{q}}}\left(X^{k}+\sigma_{k}\left(\bm{u}^{k+1}\bm{1}_{n}^{\top}+\bm{1}_{m}(\bm{v}^{k+1})^{\top}\right)\right),$ (4.5)
where
$\Psi_{k}(\bm{u},\bm{v}):=\mathcal{L}_{\sigma_{k}}(\bm{u},\bm{v},X^{k})+\frac{\tau_{k}}{2\sigma_{k}}\left\|\bm{u}-\bm{u}^{k}\right\|^{2}+\frac{\tau_{k}}{2\sigma_{k}}\left\|\bm{v}-\bm{v}^{k}\right\|^{2}.$
To truly implement ripALM for solving problem (4.2), it is essential to efficiently solve the subproblem (4.4) to find $(\bm{u}^{k+1},\bm{v}^{k+1})$ satisfying the error criterion (2.5). In the following, we shall describe how to apply a semismooth Newton (Ssn) method to achieve this goal. For simplicity, we drop the index $k$ and explicitly rewrite the subproblem (4.4) as follows:
$\hskip-5.69054pt\min\limits_{\bm{u}\in\mathbb{R}^{m},\,\bm{v}\in\mathbb{R}^{n}}\left\{\begin{aligned} &\Psi(\bm{u},\bm{v}):=-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v}+\frac{1}{2\sigma}\left\|\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right\|_{F}^{2}-\frac{1}{2\sigma}\|\bar{X}\|_{F}^{2}\\ &\quad-\mathtt{M}_{\sigma f_{\texttt{q}}}\left(\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right)+\frac{\tau}{2\sigma}\left\|\bm{u}-\bar{\bm{u}}\right\|^{2}+\frac{\tau}{2\sigma}\left\|\bm{v}-\bar{\bm{v}}\right\|^{2}\end{aligned}\right\},$ (4.6)
where $\bar{\bm{u}}$, $\bar{\bm{v}}$ and $\bar{X}$ are given. From the property of the Moreau envelope (see, for example, [2, Proposition 12.30]), we see that $\Psi$ is strongly convex and continuously differentiable with the gradient
$\nabla\Psi(\bm{u},\bm{v})=\left[\begin{matrix}\left(\mathtt{prox}_{\sigma f_{\texttt{q}}}\left(\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right)\right)\bm{1}_{n}-\bm{\alpha}+\tau\sigma^{-1}\left(\bm{u}-\bar{\bm{u}}\right)\vskip 2.84526pt\\ \left(\mathtt{prox}_{\sigma f_{\texttt{q}}}\left(\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right)\right)^{\top}\bm{1}_{m}-\bm{\beta}+\tau\sigma^{-1}\left(\bm{v}-\bar{\bm{v}}\right)\end{matrix}\right].$
Note that the proximal mapping $\mathtt{prox}_{\sigma f_{\texttt{q}}}(\cdot)$ can be easily computed as follows:
$\mathtt{prox}_{\sigma f_{\texttt{q}}}\left(Z\right)=\frac{1}{1+\lambda\sigma}\Pi_{\mathbb{R}_{+}^{m\times n}}\left(Z-\sigma{C}\right),\quad\forall\,Z\in\mathbb{R}^{m\times n}.$
Then, from the first-order optimality condition, solving problem (4.6) is equivalent to solving the following non-smooth equation:
$\nabla\Psi(\bm{u},\bm{v})=\bm{0}.$ (4.7)
In view of the nice property of $\nabla\Psi$, we are able to follow [22, 24, 25, 58] to apply a globally convergent and locally superlinearly convergent Ssn method to solve (4.7). Specifically, let $\mathcal{W}(\bm{u},\bm{v}):=\bar{X}+\sigma\left({\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-C}\right)$ and define the multifunction $\widehat{\partial}(\nabla\Psi):\mathbb{R}^{m}\times\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{(m+n)\times(m+n)}$ as follows:
[…]
where $B:=\begin{bmatrix}\bm{1}_{n}^{\top}\otimes I_{m}\\ I_{n}\otimes\bm{1}_{m}^{\top}\end{bmatrix}\in\mathbb{R}^{(m+n)\times mn}$ with $\otimes$” denoting the Kronecker product, $\texttt{vec}(\Omega)$ denotes the vectorization of $\Omega$ with $\texttt{vec}(\Omega)]_{i+(j-1)m}=\Omega_{ij}$ for any $1\leq i\leq m$ and $1\leq j\leq n$, $\mathrm{Diag}(\bm{z})$ denotes the diagonal matrix whose $i$th diagonal element is given by $z_{i}$, and $\partial{\Pi_{\mathbb{R}_{+}^{m\times n}}}(Z)$ denotes the generalized Jacobian of the Lipschitz continuous mapping $\Pi_{\mathbb{R}_{+}^{m\times n}}$ at $Z$, which is defined by
$\partial\Pi_{\mathbb{R}_{+}^{m\times n}}(Z):=\left\{\Omega\in\mathbb{R}^{m\times n}~\bigg|~\Omega_{ij}\in\left\{\begin{aligned} &\{1\},&\text{if}~~Z_{ij}>0,\\ &[0,1],&\text{if}~~Z_{ij}=0,\\ &\{0\},&\text{if}~~Z_{ij}<0,\end{aligned}\right.\quad\right\}$.
Then, using similar arguments as in the proof of [58, Proposition 4.1], one can show that $\nabla\Psi(\cdot)$ is strongly semi-smooth with respect to $\widehat{\partial}(\nabla\Psi)(\cdot)$, and thus the Ssn method is applicable. More importantly, for any $(\bm{u},\bm{v})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$, all elements of $\widehat{\partial}(\nabla\Psi)(\bm{u},\bm{v})$ are positive definite. This ensures the direct applicability of the Ssn method without the need for a specific regularity condition, such as the primal constraint nondegeneracy condition [62], highlighting the advantage of incorporating a proximal term. The detailed description of the Ssn method for solving equation (4.7) is presented in Algorithm 2. We refer readers to [22, Theorem 3.6] for its detailed convergence results.
Algorithm 2 A semismooth Newton (Ssn) method for solving equation (4.7)
Input: Choose $\bar{\mu}\in(0,1)$, $\mu\in(0,1]$, $\eta\in(0,1/2)$, ${\delta}\in(0,1)$, and an initial point $(\bm{u}^{0},\bm{v}^{0})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$. Set $t=0$.
while the termination criterion is not met do
  Step 1. Compute $\nabla\Psi(\bm{u}^{t},\bm{v}^{t})$ and choose an element $H^{t}\in\widehat{\partial}(\nabla\Psi)(\bm{u}^{t},\bm{v}^{t})$. Solve the linear system
$H^{t}\bm{d}^{t}=-\nabla\Psi(\bm{u}^{t},\bm{v}^{t}),$ (4.8)
nearly exactly by the (sparse) Cholesky decomposition with forward and backward substitutions, or approximately by the preconditioned conjugate gradient (CG) method to find $\bm{d}^{t}:=(\bm{d}_{u}^{t},\bm{d}_{v}^{t})$ such that
$\left\|H^{t}\bm{d}^{t}+\nabla\Psi(\bm{u}^{t},\bm{v}^{t})\right\|\leq\min\left\{\bar{\mu},\,\left\|\nabla\Psi(\bm{u}^{t},\bm{v}^{t})\right\|^{1+\mu}\right\}.$
  Step 2. (Line search) Find a step size $\alpha_{t}:={\delta}^{i_{t}}$, where $i_{t}$ is the smallest non-negative integer such that
$\Psi\left(\bm{u}^{t}+{\delta}^{i_{t}}\bm{d}_{u}^{t},\,\bm{v}^{t}+{\delta}^{i_{t}}\bm{d}_{v}^{t}\right)-\Psi(\bm{u}^{t},\bm{v}^{t})\leq\eta{\delta}^{i_{t}}\left\langle\nabla\Psi(\bm{u}^{t},\bm{v}^{t}),\,\bm{d}^{t}\right\rangle.$
  Step 3. Set $\bm{u}^{t+1}=\bm{u}^{t}+\alpha_{t}\bm{d}_{u}^{t}$, $\bm{v}^{t+1}=\bm{v}^{t}+\alpha_{t}\bm{d}_{v}^{t}$, $t=t+1$, and go to Step 1.
end while
4.2 Basis pursuit denoising
The basis pursuit denoising (BPDN) problem is a fundamental optimization problem in signal processing and compressed sensing, which aims to recover a sparse signal from noisy linear measurements [51]. Specifically, given a measurement $\bm{b}\in\mathbb{R}^{m}$, a dictionary $D\in\mathbb{R}^{m\times n}$, and an estimate of the noise level $\widehat{\kappa}>0$ in the data, the BPDN problem seeks a sparse solution $\bm{s}\in\mathbb{R}^{n}$ by solving the following constrained optimization problem:
$\min_{\bm{s}\in\mathbb{R}^{n}}~\|\bm{s}\|_{1}\quad\text{s.t.}\quad\|D\bm{s}-\bm{b}\|\leq\widehat{\kappa}$. (4.9)
Note that in the sparse recovery regime, one focuses on the case that $m\ll n$. By introducing an auxiliary variable $\bm{t}\in\mathbb{R}^{m}$, the above problem can be reformulated as
$\min_{\bm{s}\in\mathbb{R}^{n},\,\bm{t}\in\mathbb{R}^{m}}\;\|\bm{s}\|_{1}+\delta_{\mathcal{B}^{2}_{\widehat{\kappa}}}(\bm{t})\quad\text{s.t.}\quad D\bm{s}-\bm{t}=\bm{b},$ (4.10)
where $\mathcal{B}^{2}_{\widehat{\kappa}}\subset\mathbb{R}^{m}$ denotes the $\ell_{2}$-norm ball centered at $\bm{0}$ with radius $\widehat{\kappa}$. Then, we see that the reformulated problem (4.10) is a special case of the general problem (1.1) by letting $\bm{x}:=(\bm{s};\bm{t})\in\mathbb{R}^{n+m}$, $f(\bm{x}):=\|\bm{s}\|_{1}+\delta_{\mathcal{B}^{2}_{\widehat{\kappa}}}(\bm{t})$ and $A:=[D,-I]\in\mathbb{R}^{m\times(n+m)}$. Consequently, the proposed ripALM is directly applicable to solve the BPDN problem.
To apply ripALM, we derive the dual problem of (4.10) as follow (modulo a minus sign)
$\min_{\bm{y}\in\mathbb{R}^{m}}~~\delta_{\mathcal{B}_{1}^{\infty}}({D}^{\top}\bm{y})+\widehat{\kappa}\|\bm{y}\|-\bm{b}^{\top}\bm{y},$ (4.11)
where $\mathcal{B}_{1}^{\infty}\subset\mathbb{R}^{n}$ denotes the unit $\ell_{\infty}$-norm ball centered at $\bm{0}$. The associated augmented Lagrangian function is given by
$\mathcal{L}_{\sigma}(\bm{y},\bm{s},\bm{t})=$ $-\bm{b}^{\top}\bm{y}+\frac{1}{2\sigma}\big\|\bm{s}+\sigma{D}^{\top}\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bm{s}\|^{2}-\mathtt{M}_{\sigma\|\cdot\|_{1}}\big(\bm{s}+\sigma{D}^{\top}\bm{y}\big)$
$+\frac{1}{2\sigma}\big\|\bm{t}-\sigma\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bm{t}\|^{2}-\mathtt{M}_{\sigma\delta_{\mathcal{B}_{\widehat{\kappa}}^{2}}}\big(\bm{t}-\sigma\bm{y}\big).$
Thus, given an arbitrary initial guess $(\bm{y}^{0},\bm{s}^{0},\bm{t}^{0})\in\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{m}$, the basic iterative scheme of ripALM for solving problem (4.11) reads as follows:
$\bm{y}^{k+1}\approx\arg\min_{\bm{y}\in\mathbb{R}^{m}}\left\{\mathcal{L}_{\sigma_{k}}(\bm{y},\bm{s}^{k},\bm{t}^{k})+\frac{\tau_{k}}{2\sigma_{k}}\|\bm{y}-\bm{y}^{k}\|^{2}\right\},$ (4.12)
$\bm{s}^{k+1}=\mathtt{prox}_{\sigma_{k}\|\cdot\|_{1}}\big(\bm{s}^{k}+\sigma_{k}D^{\top}\bm{y}^{k+1}\big),$ (4.13)
$\bm{t}^{k+1}=\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}\big(\bm{t}^{k}-\sigma_{k}\bm{y}^{k+1}\big),$ (4.14)
where $\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}$ denotes the projection operator onto $\mathcal{B}_{\widehat{\kappa}}^{2}$.
To truly implement ripALM for solving problem (4.11), it is essential to efficiently solve the subproblem (4.12) to find $\bm{y}^{k+1}$ satisfying the relative error criterion (2.5). In the following, we describe how to apply the Ssn method to achieve this goal. For simplicity, we drop the index $k$ and explicitly rewrite the subproblem (4.12) as follows:
$\min_{\bm{y}\in\mathbb{R}^{m}}~~\left\{\begin{aligned} \Phi(\bm{y}):=&-\bm{b}^{\top}\bm{y}+\frac{1}{2\sigma}\big\|\bar{\bm{s}}+\sigma{D}^{\top}\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bar{\bm{s}}\|^{2}-\mathtt{M}_{\sigma\|\cdot\|_{1}}\big(\bar{\bm{s}}+\sigma{D}^{\top}\bm{y}\big)\\ &+\frac{1}{2\sigma}\big\|\bar{\bm{t}}-\sigma\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bar{\bm{t}}\|^{2}-\mathtt{M}_{\sigma\delta_{\mathcal{B}^{2}_{\widehat{\kappa}}}}\big(\bar{\bm{t}}-\sigma\bm{y}\big)+\frac{\tau}{2\sigma}\|\bm{y}-\bar{\bm{y}}\|^{2}\end{aligned}\right\}.$
where $\bar{\bm{s}}$, $\bar{\bm{t}}$ and $\bar{\bm{y}}$ are given vectors. From the property of the Moreau envelope, we see that $\Psi$ is strongly convex and continuously differentiable with the gradient
$\nabla\Phi(\bm{y})={D}\mathtt{prox}_{\sigma\|\cdot\|_{1}}\big(\bar{\bm{s}}+\sigma D^{\top}\bm{y}\big)-\Pi_{{\mathcal{B}_{\widehat{\kappa}}^{2}}}\big(\bar{\bm{t}}-\sigma\bm{y}\big)-\bm{b}+\frac{\tau}{\sigma}\big(\bm{y}-\bar{\bm{y}}\big).$ (4.15)
Then, from the first-order optimality condition, solving problem (4.12) is equivalent to solving the following non-smooth equation:
$\nabla\Phi(\bm{y})=0.$ (4.16)
Next, we define the multifunction $\widehat{\partial}(\nabla\Phi):\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ as follows:
[…]
where $\partial\mathtt{prox}_{\sigma\|\cdot\|_{1}}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n\times n}$ is defined by
[…]
and $\partial\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m\times m}$ is defined by
[…]
Using the relation $\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}(\bm{x})=\bm{x}-\mathtt{prox}_{\widehat{\kappa}\|\cdot\|_{2}}(\bm{x})$, Lemma 2.1 in [60], and similar arguments to those in [58, Proposition 4.1], one can show that $\nabla\Phi$ is strongly semismooth with respect to the multifunction $\widehat{\partial}(\nabla\Phi)$. This property enables the application of the Ssn method to solve (4.16). Since the algorithmic framework is identical to Algorithm 2, we omit the detailed description for simplicity. Moreover, note from (4.17) that, for any $\bm{y}\in\mathbb{R}^{m}$, all matrices in $\widehat{\partial}(\nabla\Phi)(\bm{y})$ are symmetric positive definite. This ensures the local superlinear (even quadratic) convergence rate of the Ssn method, without requiring additional regularity conditions. These favorable theoretical properties form the basis for the encouraging numerical performance, as observed from our numerical results in the next section.
Finally, we would like to point out that the main computational cost of the Ssn method for solving (4.16) lies in solving a sequence of linear systems of the form
$H\bm{d}=-\nabla\Phi(\bm{y}),$ (4.18)
to obtain the Newton direction $\bm{d}$, where $H\in\widehat{\partial}(\nabla\Phi)(\bm{y})$ with $\widehat{\partial}(\nabla\Phi)$ defined in (4.17). Inspired by the techniques developed in [22, Section 3.3], we can exploit the special structure of $H$ to enhance the computational efficiency of solving (4.18). However, the presence of the additional block $V$ in $H$ introduces nontrivial implementation difficulties. Detailed discussions on efficiently solving (4.18) are provided in Appendix D.
5 Numerical experiments
In this section, we conduct numerical experiments to evaluate the performance and validate the efficiency of our ripALM in Algorithm 1 for solving the QROT problem (4.1) and the BPDN problem (4.9). Specifically, we conduct experiments from the following three aspects:
  • In Section 5.1, we compare our ripALM with Snipal [24] and ciPALM [58], which represent the latest absolute-type and relative-type inexact pALMs, respectively, for solving the QROT problem on image data. These comparisons aim to illustrate how different error criteria with different tolerance parameters influence the practical numerical performance of the inexact pALM.
  • In Section 5.2, we compare our ripALM with Gurobi, Eckstein and Silva’s method [14] (denoted by ES-iALM) and a dual alternating direction method of multipliers (dADMM) presented in Appendix B, for solving large-scale QROT problems. Gurobi is one of the state-of-the-art commercial solvers, while ES-iALM is a relative-type inexact version of the vanilla ALM. Moreover, dADMM is a widely used first-order method for large-scale constrained convex optimization; see, for example, [6, 17]. In addition, it is worth noting that many other first-order methods can also be applied to the QROT problem, such as the Chambolle–Pock algorithm [8], the Condat–Vũ algorithm [9, 52], and the linearized ALM [53, 55]. However, our preliminary tests show that these methods often perform comparably to, or worse than, dADMM. For this reason, and for the sake of simplicity, we omit them from our comparisons and only select the dADMM in Appendix B as a representative first-order method.
  • In Section 5.3, we compare our ripALM with SPGL1 [51], Mosek, and the dADMM presented in Appendix E, for solving the BPDN problem. SPGL1 is a state-of-the-art solver specifically designed for BPDN, while Mosek is a highly efficient commercial solver that can be applied to solve the second-order cone programming reformulation of the BPDN problem; see Appendix F for more details.
All experiments are run in Matlab R2023a on a PC with Intel processor i7-12700K@3.60GHz (with 12 cores and 20 threads) and 64GB of RAM, equipped with a Windows OS. The implementation details are given as follows.
Termination conditions. We first describe the termination conditions for QROT. Let $\mathcal{Z}(\bm{u},\bm{v},X):=C+\lambda X-\bm{u}\bm{1}_{n}^{\top}-\bm{1}_{m}\bm{v}^{\top}$. The Karush-Kuhn-Tucker (KKT) system for problem (4.1) and its dual (4.3) is given by
$\begin{array}[]{l}X\bm{1}_{n}=\bm{\alpha},~~X^{\top}\bm{1}_{m}=\bm{\beta},~~\langle X,\,\mathcal{Z}(\bm{u},\bm{v},X)\rangle=0,~~X\geq 0,~~\mathcal{Z}(\bm{u},\bm{v},X)\geq 0.\end{array}$ (5.1)
Note that $(\bm{u},\bm{v},X)$ satisfies the KKT system (5.1) if and only if $X$ solves (4.1) and $(\bm{u},\bm{v})$ solves (4.3), respectively. Based on (5.1), we define the relative KKT residual for any $(X,\,\bm{u},\,\bm{v})$ as follows:
$\Delta_{\rm kkt}^{\rm qrot}(\bm{u},\bm{v},X):=\max\left\{\Delta_{p}^{\rm qrot}(X),\,\Delta_{d}^{\rm qrot}(\bm{u},\bm{v},X),\,\Delta_{c}^{\rm qrot}(\bm{u},\bm{v},X)\right\},$
where
$\;\Delta_{p}^{\rm qrot}(X):=\max\left\{\frac{\|X\bm{1}_{n}-\bm{\alpha}\|}{1+\|\bm{\alpha}\|},\,\frac{\|X^{\top}\bm{1}_{m}-\bm{\beta}\|}{1+\|\bm{\beta}\|},\,\frac{\|\min\{X,\,0\}\|_{F}}{1+\|X\|_{F}}\right\},$
$\;\Delta_{d}^{\rm qrot}(\bm{u},\bm{v},X):=\frac{\|\min\{\mathcal{Z}(\bm{u},\,\bm{v},\,X),\,0\}\|_{F}}{1+\|C\|_{F}},\quad\;\Delta_{c}^{\rm qrot}(\bm{u},\bm{v},X):=\frac{\left|\langle X,\,\mathcal{Z}(\bm{u},\,\bm{v},\,X)\rangle\right|}{1+\|C\|_{F}}.$
Moreover, we define the relative duality gap as follows:
$\Delta_{\rm gap}^{\rm qrot}(\bm{u},\,\bm{v},\,X):=\cfrac{\left|\mathtt{pobj}(X)-\mathtt{dobj}(\bm{u},\bm{v})\right|}{1+\left|\mathtt{pobj}(X)\right|+\left|\mathtt{dobj}(\bm{u},\bm{v})\right|},$
where $\mathtt{pobj}(X):=f_{\texttt{q}}(X)$ and $\mathtt{dobj}(\bm{u},\bm{v}):=-f_{\texttt{q}}^{*}\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)+\bm{\alpha}^{\top}\bm{u}+\bm{\beta}^{\top}\bm{v}$. Using these relative residuals, we will terminate ripALM when it returns a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ satisfying
$\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k}):=\max\left\{\Delta_{\rm kkt}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k}),\,\Delta_{\rm gap}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})\right\}<10^{-6}.$
Similarly, we set up the termination conditions for BPDN based on the associated KKT system. Specifically, given an output solution $(\bm{s},\bm{t},\bm{y})$, we define
$\Delta_{p}^{\rm bpdn}(\bm{s},\bm{t})$ $:=\cfrac{\left\|D\bm{s}-\bm{b}-\bm{t}\right\|}{1+\|\bm{b}\|},$
$\Delta_{d}^{\rm bpdn}(\bm{s},\bm{t},\bm{y})$ $:=\cfrac{\sqrt{\big\|\bm{s}-\mathtt{prox}_{\|\cdot\|_{1}}\left(\bm{s}+{D}^{\top}\bm{y}\right)\big\|^{2}+\big\|\bm{t}-\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}\left(\bm{t}-\bm{y}\right)\big\|^{2}}}{1+\|D\|_{F}},$
${\Delta}_{\rm gap}^{\rm bpdn}(\bm{s},\bm{t},\bm{y})$ $:=\cfrac{\left|\mathtt{pobj}(\bm{s},\bm{t})-\mathtt{dobj}(\bm{y})\right|}{1+\left|\mathtt{pobj}(\bm{s},\bm{t})\right|+\left|\mathtt{dobj}(\bm{y})\right|},$
where $\mathtt{pobj}(\bm{s},\bm{t}):=\|\bm{s}\|_{1}$ and $\mathtt{dobj}(\bm{y}):=-\widehat{\kappa}\|\bm{y}\|+\bm{b}^{\top}\bm{y}$. Then, we terminate ripALM when it returns $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ such that
$\Delta_{\rm res}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k}):=\max\left\{\Delta_{p}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k}),\,\Delta_{d}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k}),\,\Delta_{\rm gap}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})\right\}<10^{-6}.$
Initial points. For the experiments in Section 5.1, we simply initialize all inexact pALMs at the origin to amplify the impact of inexactness. For the experiments in Section 5.2 and Section 5.3, we will employ a warm-start strategy for more efficiency. Indeed, our numerical experience (see, e.g., [24, 26, 58]) have suggested that a reasonably good initial point would benefit the practical performance of such a “pALM + Ssn” algorithmic framework. Therefore, for the experiments in Section 5.2, we first employ an inexact Bregman proximal gradient method (iBPGM) with Sinkhorn’s algorithm as a subsolver for solving the QROT problem (4.1) to generate an initial point for our ripALM; see Appendix C for more details. Specifically, we terminate this iBPGM as long as it produces a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ such that $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})<10^{-3}$, or its iteration number reaches 500. Moreover, we also apply this warm-start strategy to ES-iALM in Section 5.2 to ensure a fair comparison. For the experiments in Section 5.3, we use dADMM in Appendix E to generate an initial point. Specifically, we terminate dADMM as long as it produces a point $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ such that $\Delta_{\mathrm{res}}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})<10^{-3}$, or its iteration number reaches 100. Note that the time consumed during the above warm-starting phase is included in the total computational time for both ripALM and ES-iALM. In addition, the initial error variable $\bm{w}^{0}$ in ripALM and ES-iALM is always set to $\bm{0}$.
For the Ssn method in Algorithm 2, we will initialize it with the origin at the first pALM iteration and then employ a warm-start strategy thereafter. Specifically, at each pALM iteration, we initialize the Ssn method with the approximate solution obtained by the Ssn method in the previous pALM iteration.
Hyperparameters. Our ripALM as well as Snipal and ciPALM require appropriate choices of $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ to achieve superior performance. In our experiments, for all these algorithms, we simply set $\sigma_{k}=\min\big\{10^{4},\,\max\big\{10^{-4},\,1.5^{k}\big\}\big\}$ and $\tau_{k}\equiv 5$ for all $k\geq 0$, without employing any sophisticated parameter tuning. These choices of $\{\tau_{k}\}$ and $\{\sigma_{k}\}$ can satisfy the conditions required in Theorems 3.1 and 3.2; see also Remark 3.1. Moreover, we would like to emphasize that adopting more refined updating rules for $\sigma_{k}$ and $\tau_{k}$ is certainly possible and may further improve the numerical performance. In addition, for the Ssn method in Algorithm 2, we set $\eta=10^{-4}$, $\delta=0.5$, $\bar{\mu}=10^{-3}$ and $\mu=0.2$.
5.1 Comparison with SNIPAL and ciPALM
In this part of experiments, we compare our ripALM with Snipal [24] and ciPALM [58] for solving the QROT problem (4.1) to illustrate how different error criteria with different tolerance parameters influence the numerical performance of the inexact pALM. Note that both ripALM and ciPALM are of relative-type and only require a single tolerance parameter $\rho$, as shown in (2.5) and (2.8), respectively, while Snipal is of absolute-type and requires two summable tolerance parameter sequences $\{\varepsilon_{k}\}$ and $\{\delta_{k}\}$, as shown in (2.7). For simplicity, in our comparisons, we set $\varepsilon_{k}=\varepsilon_{0}/(k+1)^{p},\;\delta_{k}=\delta_{0}/(k+1)^{q}$ with $\varepsilon_{0}=\delta_{0}\in\big\{0.01,1\big\}$ and $p,q\in\big\{1.1,\,2.1,\,3.1\big\}$ (hence, there are 18 combinations in total) for Snipal. For ripALM and ciPALM, we consider $\rho\in\left\{0.01,0.05,0.1,0.3,0.5,0.7,0.9,0.99,0.999\right\}$ (hence, there are 9 choices).
We use images from the ClassicImages and MicroscopyImages classes in the DOTmark collection [44], which serves as a benchmark dataset for the OT problem and its variants, to generate QROT instances. Note that the images in the ClassicImages (or MicroscopyImages) class consist of 10 different images, each with different resolutions of $32\times 32$, $64\times 64$, $128\times 128$ and $512\times 512$. Thus, for each resolution, we can pair any two different images within a class and compute the QROT problem, resulting in 90 QROT problems (45 from ClassicImages and 45 from MicroscopyImages). Moreover, the cost matrix $C$ is obtained by calculating the squared Euclidean distances between pixels.
Tables 2 and 3 present the average results (over 90 instances) for $64\times 64$ resolution with the regularization parameter $\lambda\in\{1,\,0.1\}$. From the results, we see that the performance of all algorithms depends on the choices of tolerance parameters, and with proper tuning, their performances can be comparable. This is indeed expected because all the algorithms essentially employ the same “pALM + Ssn” framework and differ only in their error criteria for solving subproblems. Notably, since ripALM and ciPALM involve only a single tolerance parameter $\rho\in[0,1)$, they would be more user-friendly and easier to tune through a simple one-dimensional grid search. This supports the main motivation for developing a relative-type error criterion. Moreover, one can see that ripALM always outperforms ciPALM in terms of the total number of Ssn iterations and exhibits a more robust performance, particularly when a larger $\rho$ is chosen, leading to a noticeable reduction in computational time. This difference may be attributed to the extra correction step required by ciPALM to guarantee convergence, which tends to considerably increase the number of Ssn iterations [^4] One possible reason is that the variable obtained after the correction step may no longer serve as a suitable initial point for the Ssn method in the next iteration. needed to satisfy its error criterion (2.8), even though these extra iterations do not contribute to the progress of the pALM iterations. These observations further support the advantage of the vanilla inexact pALM framework advocated in this paper. As shown in Tables 2 and 3, our ripALM clearly exhibits greater robustness and efficiency. In addition, its performance remains relatively stable across different values of $\rho$, with slight advantage for $\rho\geq 0.5$.
Table 2: Comparisons between ripALM, Snipal and ciPALM under different choices of tolerance parameters for solving the QROT problem (4.1), where $\lambda=1$ and the instances are generated using images with the resolution of $64\times 64$ in the ClassicImages and MicroscopyImages classes from the DOTmark collection. In the table, $\Delta_{\mathrm{res}}^{\rm qrot}$” denotes the terminating $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$; “#” denotes the number of iterations (the total number of Ssn iterations is given in the bracket); “time” denotes the computational time in seconds.
Snipal
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
(1, 1.1, 1.1) 7.48e-07 17 (51) 8.90 (0.01, 1.1, 1.1) 7.48e-07 17 (67) 11.01
(1, 1.1, 2.1) 7.48e-07 17 (58) 9.76 (0.01, 1.1, 2.1) 7.48e-07 17 (70) 11.42
(1, 1.1, 3.1) 7.48e-07 17 (62) 10.23 (0.01, 1.1, 3.1) 7.48e-07 17 (72) 11.65
(1, 2.1, 1.1) 7.48e-07 17 (51) 8.89 (0.01, 2.1, 1.1) 7.48e-07 17 (67) 11.01
(1, 2.1, 2.1) 7.48e-07 17 (58) 9.74 (0.01, 2.1, 2.1) 7.48e-07 17 (70) 11.43
(1, 2.1, 3.1) 7.48e-07 17 (62) 10.21 (0.01, 2.1, 3.1) 7.48e-07 17 (72) 11.66
(1, 3.1, 1.1) 7.48e-07 17 (51) 8.89 (0.01, 3.1, 1.1) 7.48e-07 17 (67) 11.01
(1, 3.1, 2.1) 7.48e-07 17 (58) 9.75 (0.01, 3.1, 2.1) 7.48e-07 17 (70) 11.41
(1, 3.1, 3.1) 7.48e-07 17 (62) 10.22 (0.01, 3.1, 3.1) 7.48e-07 17 (72) 11.65
ripALM ciPALM
$\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
0.999 7.48e-07 17 (42) 8.01 0.999 7.48e-07 17 (164) 71.95
0.99 7.48e-07 17 (42) 8.02 0.99 7.55e-07 17 (163) 71.53
0.9 7.48e-07 17 (42) 8.03 0.9 7.40e-07 17 (156) 67.33
0.7 7.48e-07 17 (42) 8.10 0.7 7.48e-07 17 (142) 58.87
0.5 7.48e-07 17 (43) 8.18 0.5 7.48e-07 17 (125) 48.21
0.3 7.48e-07 17 (44) 8.32 0.3 7.48e-07 17 (105) 36.16
0.1 7.48e-07 17 (47) 8.68 0.1 7.48e-07 17 (79) 20.00
0.05 7.48e-07 17 (49) 8.93 0.05 7.48e-07 17 (70) 14.73
0.01 7.48e-07 17 (55) 9.59 0.01 7.48e-07 17 (65) 11.30
Table 3: Same as Table 2 but for $\lambda=0.1$.
Snipal
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
(1, 1.1, 1.1) 6.19e-07 20 (92) 27.72 (0.01, 1.1, 1.1) 6.19e-07 20 (110) 29.93
(1, 1.1, 2.1) 6.19e-07 20 (100) 28.60 (0.01, 1.1, 2.1) 6.19e-07 20 (113) 30.32
(1, 1.1, 3.1) 6.19e-07 20 (104) 29.11 (0.01, 1.1, 3.1) 6.19e-07 20 (115) 30.56
(1, 2.1, 1.1) 6.19e-07 20 (92) 27.74 (0.01, 2.1, 1.1) 6.19e-07 20 (110) 29.93
(1, 2.1, 2.1) 6.19e-07 20 (100) 28.63 (0.01, 2.1, 2.1) 6.19e-07 20 (113) 30.32
(1, 2.1, 3.1) 6.19e-07 20 (104) 29.10 (0.01, 2.1, 3.1) 6.19e-07 20 (115) 30.56
(1, 3.1, 1.1) 6.19e-07 20 (92) 27.72 (0.01, 3.1, 1.1) 6.19e-07 20 (110) 29.93
(1, 3.1, 2.1) 6.19e-07 20 (100) 28.60 (0.01, 3.1, 2.1) 6.19e-07 20 (113) 30.32
(1, 3.1, 3.1) 6.19e-07 20 (104) 29.09 (0.01, 3.1, 3.1) 6.19e-07 20 (115) 30.57
ripALM ciPALM
$\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\rho$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
0.999 6.20e-07 20 (81) 26.68 0.999 6.19e-07 20 (230) 111.70
0.99 6.20e-07 20 (81) 26.68 0.99 6.19e-07 20 (232) 113.24
0.9 6.20e-07 20 (82) 26.72 0.9 6.19e-07 20 (222) 105.97
0.7 6.20e-07 20 (82) 26.78 0.7 6.19e-07 20 (200) 91.75
0.5 6.20e-07 20 (83) 26.85 0.5 6.19e-07 20 (181) 79.50
0.3 6.19e-07 20 (85) 27.00 0.3 6.19e-07 20 (156) 63.09
0.1 6.19e-07 20 (88) 27.39 0.1 6.19e-07 20 (122) 41.27
0.05 6.19e-07 20 (90) 27.68 0.05 6.19e-07 20 (115) 36.27
0.01 6.19e-07 20 (96) 28.39 0.01 6.19e-07 20 (108) 31.14
5.2 Comparison with Gurobi, dADMM and ES-iALM
In this part of experiments, we compare our ripALM with Gurobi (version 12.0.1), dADMM and ES-iALM for solving large-scale QROT problems. For our ripALM, we set $\rho=0.99$ based on the numerical observations from the previous section. For ES-iALM, we adopt the same settings for $\rho$, $\sigma_{k}$}, and the termination conditions as those used in ripALM. Moreover, we employ FISTA [3] to solve the ALM-subproblems, with the maximum number of total inner iterations set to 15000.555Note that since ES-iALM does not involve a proximal term, the Ssn method may not be applicable to its subproblems, as the nonsingularity of the generalized Jacobian cannot be guaranteed. Therefore, we employ FISTA as the subsolver in ES-iALM. For Gurobi, we use its default termination conditions and set the corresponding termination tolerances as $10^{-6}$, aligning it with our tolerance. For dADMM, we initialize the penalty parameter as $\sigma_{0}=0.01\|C\|_{F}^{-1}$, and dynamically adjust it based on the primal-dual residuals, following the approach described in [57, Section 5.1], to achieve superior empirical performance. Moreover, we will terminate dADMM when it returns a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})<10^{-6}$, or its number of iterations reaches 10000.
We follow [59, Section 5.1] to generate a random instance. Specifically, we first generate two discrete probability distributions $\big\{(a_{i},\bm{p}_{i})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:i=1,\cdots,m\big\}$ and $\big\{(b_{j},\bm{q}_{j})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:j=1,\cdots,n\big\}$. Here, $\bm{a}:=(a_{1},\cdots,a_{m})^{\top}$ and $\bm{b}:=(b_{1},\cdots,b_{n})^{\top}$ are probabilities/weights, which are generated from the uniform distribution on the open interval $(0,1)$ and further normalized such that $\sum^{m}_{i}a_{i}=\sum^{n}_{j}b_{j}=1$. Moreover, $\{\bm{p}_{i}\}$ and $\{\bm{q}_{j}\}$ are support points whose entries are drawn from a five-component multivariate Gaussian mixture distribution, with a mean vector $(-20,10,0,10,20)^{\top}$ and a variance vector $(5,5,5,5,5)^{\top}$, using randomly generated weights. Then, the cost matrix $C$ is generated by $c_{ij}=\|\bm{p}_{i}-\bm{q}_{j}\|^{2}$ for $1\leq i\leq m$ and $1\leq j\leq n$ and normalized by dividing (element-wise) by its maximal entry.
We then generate a set of random instances with $m=n\in\{1000,\,2000,\dots,10000\}$. For each $m$, we generate 10 instances with different random seeds, and present the average numerical performances of Gurobi, dADMM, ES-iALM, and ripALM in Table 4 and Figure 1, with $\lambda\in\{1,\,0.1\}$. We observe that both ripALM and Gurobi are able to solve the tested problems accurately, in the sense that the residual $\Delta_{\mathrm{res}}^{\rm qrot}$ falls below $10^{-6}$. In contrast, ES-iALM often fails to attain solutions of comparable accuracy, even after reaching its maximum iteration limit and incurring substantially higher computational costs. This is mainly due to the limited efficiency of the first-order subsolver used in ES-iALM. Moreover, dADMM performs even worse: despite consuming a similar amount of computational time as ES-iALM, it yields solutions of considerably lower quality. In addition, we have also observed that Gurobi can be rather time-consuming and memory-consuming for large-scale problems. As an example, for the case where $m=n=10000$, a large-scale QP containing $10^{8}$ nonnegative variables and 20000 equality constraints was solved. One can see that Gurobi is around $2\!\sim\!4$ times slower than our ripALM. In addition, Gurobi may lack robustness, especially for solving large-scale problems. Indeed, as observed from Figure 1, the computational times taken by Gurobi vary a lot among 10 randomly generated instances. In conclusion, our ripALM shows potential for greater efficiency and robustness in solving large-scale QROT problems.
Table 4: Numerical results of Gurobi, dADMM, ES-iALM, and ripALM for solving the QROT problem (4.1). In the table, $\Delta_{\mathrm{res}}^{\rm qrot}$” denotes the terminating $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$; “#” denotes the number of iterations (the total number of Ssn/FISTA iterations is given in the bracket); “time” denotes the computational time.
Gurobi dADMM ES-iALM ripALM
$m=n$ $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\Delta_{\mathrm{res}}^{\rm qrot}$ # time $\Delta_{\mathrm{res}}^{\rm qrot}$ # time
$\lambda=1$
1000 2.78e-07 33 4.70 2.91e-06 10000 15.64 6.31e-07 16 (11483) 15.11 5.21e-07 16 (38) 1.71
2000 4.84e-07 36 21.82 1.02e-05 10000 101.69 2.23e-06 16 (14605) 105.80 7.16e-07 16 (50) 7.91
3000 4.58e-07 39 56.69 2.24e-05 10000 244.25 2.65e-05 16 (15000) 256.41 5.90e-07 17 (59) 19.35
4000 3.44e-07 40 112.30 5.61e-05 10000 440.24 1.11e-05 15 (15000) 460.29 6.48e-07 17 (63) 36.82
5000 4.50e-07 44 191.43 7.21e-05 10000 688.22 5.85e-06 15 (15000) 720.98 6.31e-07 17 (68) 61.49
6000 5.80e-07 45 290.88 9.79e-05 10000 992.49 1.37e-05 15 (15000) 1034.33 6.35e-07 17 (71) 89.46
7000 4.71e-07 47 441.00 1.48e-04 10000 1349.22 1.58e-05 15 (15000) 1404.71 4.98e-07 18 (77) 129.41
8000 5.78e-07 47 625.84 1.87e-04 10000 1762.69 2.51e-05 14 (15000) 1830.78 4.55e-07 18 (81) 175.42
9000 4.70e-07 50 774.31 1.94e-04 10000 2229.83 2.17e-05 14 (15000) 2323.49 5.64e-07 18 (82) 225.48
10000 6.90e-07 50 998.24 2.09e-04 10000 2757.73 3.28e-05 14 (15000) 2864.61 5.04e-07 18 (85) 285.40
$\lambda=0.1$
1000 3.89e-07 32 4.63 3.21e-05 10000 16.11 8.33e-05 16 (15000) 18.87 6.44e-07 18 (67) 1.93
2000 2.71e-07 36 21.88 1.26e-04 10000 102.04 7.87e-05 15 (15000) 107.03 6.49e-07 18 (77) 9.69
3000 3.88e-07 40 57.76 3.10e-04 10000 243.63 8.38e-05 15 (15000) 255.45 6.73e-07 19 (86) 27.50
4000 3.81e-07 41 120.33 5.90e-04 10000 439.32 6.54e-05 14 (15000) 461.86 6.81e-07 19 (93) 56.32
5000 5.93e-07 47 201.37 7.22e-04 10000 689.08 3.25e-05 15 (15000) 729.19 6.18e-07 19 (99) 97.73
6000 5.47e-07 48 308.46 1.12e-03 10000 994.07 3.19e-04 14 (15000) 1047.92 6.50e-07 20 (110) 155.45
7000 4.46e-07 46 409.37 1.40e-03 10000 1354.34 1.67e-04 14 (15000) 1420.33 6.10e-07 20 (116) 225.95
8000 1.37e-06 47 615.47 1.80e-03 10000 1762.95 3.67e-04 14 (15000) 1852.26 6.25e-07 20 (120) 306.10
9000 8.38e-07 50 779.09 2.04e-03 10000 2230.05 1.84e-04 14 (15000) 2343.77 6.82e-07 20 (127) 408.59
10000 7.11e-07 52 1051.46 1.99e-03 10000 2750.64 1.44e-04 14 (15000) 2897.32 7.40e-07 20 (128) 504.35
Figure 1: Comparisons among Gurobi, dADMM, ES-iALM, and ripALM for solving the QROT problem (4.1) with $m=n\in\{1000,\,2000,\cdots,10000\}$. The shaded region indicates the maximum and minimum computational times taken to solve 10 instances for each problem dimension.
5.3 Experiments for BPDN
In this section, we compare our ripALM with Mosek (version 11.0.27), SPGL1 (version 2.1), and dADMM (see Appendix E) for solving the BPDN problem (4.9). For ripALM, we set $\rho=0.99$ as in the previous section. For dADMM, we initialize the penalty parameter as $\sigma=1$, and dynamically update it based on the primal-dual residuals, following the strategy in [57, Section 5.1]. We terminate dADMM when it returns a point $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})<10^{-6}$ or reaches 10000 iterations. For SPGL1, we use its default termination conditions, set the tolerance as $10^{-6}$, (aligning it with our tolerance), and set the maximum number of iterations as 10000. For Mosek, we use its default termination conditions, but adopt two tolerance levels: a tolerance of $10^{-6}$ (denoted as Mosek (1e-6)) to match our tolerance and a tighter tolerance of $10^{-10}$ (denoted as Mosek (1e-10)) to generate a high-precision solution to use as a benchmark.
We construct the tested BPDN problems using both synthetic datasets and UCI datasets [^6] Available at: https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/regression.html. The synthetic datasets are generated as follows. Given a triple $(m,n,s)$, we first generate a dictionary matrix $D\in\mathbb{R}^{m\times n}$ with i.i.d. standard Gaussian entries. We then uniformly select an index set $\mathcal{S}\subset\{1,\cdots,n\}$ with cardinality $|\mathcal{S}|=s$ and construct an $s$-sparse signal $\bar{\bm{s}}\in\mathbb{R}^{n}$, whose entries on $\mathcal{S}$ are drawn independently from the standard Gaussian distribution and are zero elsewhere. The measurement vector is then formed as $\bm{b}=D\bar{\bm{s}}+\delta\bm{\zeta}$, where $\bm{\zeta}\in\mathbb{R}^{m}$ is a noise vector with i.i.d. standard Gaussian entries and $\delta>0$ is the noise level. For UCI datasets, we choose housing7 and mpg7 to obtain $(D,\bm{b})$, where $D$ has dimensions $506\times 77520$ for housing7 and $392\times 3432$ for mpg7. Note that housing7 and mpg7 are expanded versions of their datasets, with the suffix ‘7’ indicating that an order $7$ polynomial is used to generate the basis functions; see [22, Section 4.1] for more details. Finally, to evaluate the solution quality, we compute the relative feasibility residual and normalized objective gap, defined respectively as:
$\texttt{feas}(\bm{s}):=\frac{\max\left\{\|D\bm{s}-\bm{b}\|_{2}-\widehat{\kappa},\,0\right\}}{1+\|\bm{b}\|_{2}},\quad\text{ and }\quad\texttt{nobj}(\bm{s}):=\frac{\big|\|\bm{s}\|_{1}-\|\bm{s}^{*}\|_{1}\big|}{1+\|\bm{s}^{*}\|_{1}},$
where $\bm{s}^{*}$ denote the solution returned by Mosek (1e-10).
Table 5 reports the average computational results for (m,n,s)=(0.1i,i, 0.02i)(m,n,s)=(0.1i,\,i,\,0.02i) with $i\in\{10000,20000,30000\}$ over $10$ randomly generated BPDN instances. In these experiments, the parameter $\widehat{\kappa}$ in (4.9) is set to $\widehat{\kappa}=\delta\left\|\bm{\zeta}\right\|_{2}$ with $\delta\in\{0.01,\,0.1\}$. The results demonstrate the superior performance of ripALM: it consistently achieves the smallest normalized objective gap, while requiring the least computational time and maintaining favorable feasibility residuals. Table 6 reports the results on two UCI data sets, where we choose $\widehat{\kappa}=\delta\|\bm{b}\|$ with $\delta\in\{0.01,\,0.1\}$. These results further show the encouraging performance of our ripALM on the real-world datasets.
Table 5: Numerical results of Mosek, SPGL1, dADMM, and ripALM for solving the BPDN problem (4.9) with $(m,n,s)=(0.1i,i,0.02i)$ and $\widehat{\kappa}=\delta\left\|\bm{\zeta}\right\|_{2}$. In the table, “nobj” denotes the normalized objective gap; “feas” denotes the relative feasibility residual; “#” denotes the number of iterations (the total number of Ssn iterations is given in the bracket); “time” denotes the computational time.
$\delta=0.01$ $\delta=0.1$
$i$ method nobj feas # time nobj feas # time
10000 Mosek (1e-10) 0 7.94e-12 26 6.80 0 1.04e-11 24 6.31
Mosek (1e-6) 1.09e-05 3.42e-07 19 5.52 1.00e-05 3.12e-07 18 5.21
SPGL1 4.72e-05 0 4520 28.32 1.35e-05 0 703 6.15
dADMM 7.49e-06 4.49e-12 10000 71.65 1.95e-06 8.98e-13 4081 31.20
ripALM 1.35e-08 1.31e-09 9 (58) 2.26 4.35e-09 2.89e-09 9 (46) 1.87
20000 Mosek (1e-10) 0 7.50e-12 32 36.20 0 8.51e-12 28 31.61
Mosek (1e-6) 1.12e-05 3.42e-07 21 25.79 1.05e-05 3.20e-07 20 25.02
SPGL1 4.83e-05 0 7991 219.83 3.71e-05 0 997 34.68
dADMM 8.75e-06 1.83e-12 10000 286.40 1.93e-06 5.38e-13 5391 179.74
ripALM 2.83e-09 7.64e-10 9 (81) 12.63 1.06e-08 3.47e-10 10 (62) 10.17
30000 Mosek (1e-10) 0 9.16e-12 30 83.94 0 7.44e-12 32 87.60
Mosek (1e-6) 8.35e-06 2.51e-07 22 67.40 1.31e-05 3.95e-07 21 64.39
SPGL1 5.28e-05 0 8156 479.54 1.69e-05 0 1038 80.32
dADMM 9.36e-06 2.06e-12 10000 601.85 1.88e-06 2.64e-13 6121 453.56
ripALM 5.28e-09 7.82e-10 8 (81) 32.52 4.49e-08 3.88e-10 9 (71) 28.50
Table 6: Same as Table 5 but for the UCI datasets with $\widehat{\kappa}=\delta\|\bm{b}\|$.
$\delta=0.01$ $\delta=0.1$
datasets method nobj feas # time nobj feas # time
housing7 Mosek (1e-10) 0 2.37e-11 36 44.94 0 2.40e-12 27 35.05
Mosek (1e-6) 1.83e-05 4.60e-08 22 25.42 4.31e-05 1.44e-08 19 23.89
SPGL1 1.76e-01 1.42e-02 10000 254.94 9.21e-03 0 1539 51.72
dADMM 4.98e-06 0 10000 237.17 1.74e-06 0 3201 76.53
ripALM 6.58e-06 0 22 (74) 2.28 2.08e-09 0 18 (45) 1.15
mpg7 Mosek (1e-10) 0 1.49e-09 26 1.03 0 1.52e-13 16 0.65
Mosek (1e-6) 1.45e-07 2.98e-08 17 0.62 8.92e-06 1.82e-08 13 0.53
SPGL1 8.82e-01 3.85e-02 10000 3.00 5.35e-03 0 443 0.22
dADMM 1.48e-04 0 10000 3.05 2.33e-06 0 2001 0.60
ripALM 1.03e-06 1.73e-08 30 (126) 0.49 2.06e-07 1.58e-10 18 (23) 0.02
6 Conclusions
In this paper, we developed a relative-type inexact proximal augmented Lagrangian method (ripALM) for solving a class of linearly constrained convex optimization problems. The proposed ripALM is the first relative-type inexact version of the vanilla pALM with provable convergence guarantees. By employing a suitable relative-type error criterion, it simplifies implementation and parameter tuning, compared to the absolute-type inexact pALM. We conducted a thorough convergence analysis and demonstrated the competitive efficiency of our ripALM through numerical experiments on solving quadratically regularized optimal transport problems and basis pursuit denoising problems. Since our model covers a wide range of applications, future work may explore the potential of applying ripALM to other practical problems, including general regularized optimal transport problems as studied in [58]. Additionally, it would be interesting to extend ripALM to a broader class of convex optimization problems, thereby further enhancing its applicability.
Acknowledgements
We thank the associate editor and two anonymous reviewers for their valuable comments and suggestions, which have helped to improve the quality of this paper. The research of Lei Yang is supported in part by the National Key Research and Development Program of China under grant 2023YFB3001704 and the National Natural Science Foundation of China under grant 12301411. The research of Kim-Chuan Toh is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 grant (MOE-T2EP20224-0017).
Appendix A Missing Proofs in Section 3
A.1 Proof of Theorem 3.1
Proof.
Statement (i). First, recall that $\ell(\bm{y},\bm{x})=-\bm{b}^{\top}\bm{y}+\langle\bm{x},A^{\top}\bm{y}\rangle-f(\bm{x})$, which is convex in $\bm{y}$ and concave in $\bm{x}$. We see that $\partial\ell(\bm{y},\bm{x})=\left\{A\bm{x}-\bm{b}\right\}\times\left\{v-A^{\top}\bm{y}\mid v\in\partial f(\bm{x})\right\}$. By using the relations in (2.5) and (2.6), along with some manipulations, we can obtain the following results for all $k\geq 0$:
$\big(\Delta^{k+1}-\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k+1}-\bm{y}^{k}),\,\sigma_{k}^{-1}(\bm{x}^{k}-\bm{x}^{k+1})\big)\in\partial\ell(\bm{y}^{k+1},\bm{x}^{k+1}),$ (A.1)
$2\big|\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle\big|+\big\|\sigma_{k}\Delta^{k+1}\big\|^{2}\leq\rho\big(\big\|\bm{x}^{k+1}-\bm{x}^{k}\big\|^{2}+\tau_{k}\big\|\bm{y}^{k+1}-\bm{y}^{k}\big\|^{2}\big),$ (A.2)
$\bm{w}^{k+1}=\bm{w}^{k}-\sigma_{k}\Delta^{k+1}.$ (A.3)
Let $(\bm{y}^{*},\bm{x}^{*})\in\mathbb{R}^{M}\times\mathbb{R}^{N}$ be an arbitrary saddle point of $\ell$ and hence $(\bm{0},\bm{0})\in\partial\ell(\bm{y}^{*},\bm{x}^{*})$. For all $k\geq 0$,
$\|\bm{x}^{k}-\bm{x}^{*}\|^{2}$ $=\|\bm{x}^{k}-\bm{x}^{k+1}+\bm{x}^{k+1}-\bm{x}^{*}\|^{2}$
$=\|\bm{x}^{k}-\bm{x}^{k+1}\|^{2}+2\langle\bm{x}^{k}-\bm{x}^{k+1},\,\bm{x}^{k+1}-\bm{x}^{*}\rangle+\|\bm{x}^{k+1}-\bm{x}^{*}\|^{2}.$
By letting $\bm{\xi}^{k+1}:=\sigma_{k}^{-1}(\bm{x}^{k}-\bm{x}^{k+1})$, the above equation can be reformulated as
$\|\bm{x}^{k+1}-\bm{x}^{*}\|^{2}=\|\bm{x}^{k}-\bm{x}^{*}\|^{2}-2\sigma_{k}\langle\bm{x}^{k+1}-\bm{x}^{*},\,\bm{\xi}^{k+1}\rangle-\|\bm{x}^{k+1}-\bm{x}^{k}\|^{2}.$ (A.4)
Then, using the relation $\bm{w}^{k+1}=\bm{w}^{k}-\sigma_{k}\Delta^{k+1}$ (by (A.3)), we see that
$\|\bm{w}^{k+1}-\bm{y}^{*}\|^{2}=$ $\;\|\bm{w}^{k}-\sigma_{k}\Delta^{k+1}-\bm{y}^{*}\|^{2}$ (A.5)
$=$ $\;\|\bm{w}^{k}-\bm{y}^{*}\|^{2}-2\langle\bm{w}^{k}-\bm{y}^{*},\,\sigma_{k}\Delta^{k+1}\rangle+\|\sigma_{k}\Delta^{k+1}\|^{2}$
$=$ $\;\|\bm{w}^{k}-\bm{y}^{*}\|^{2}-2\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle+\|\sigma_{k}\Delta^{k+1}\|^{2}$
$\;-2\sigma_{k}\langle\bm{y}^{k+1}-\bm{y}^{*},\,\bm{\theta}^{k+1}\rangle-2\tau_{k}\langle\bm{y}^{k+1}-\bm{y}^{*},\,\bm{y}^{k+1}-\bm{y}^{k}\rangle.$
where $\bm{\theta}^{k+1}:=\Delta^{k+1}-\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k+1}-\bm{y}^{k})$. Similarly,
$\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{*}\|^{2}=\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}-\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{k}\|^{2}+2\tau_{k}\langle\bm{y}^{k+1}-\bm{y}^{*},\,\bm{y}^{k+1}-\bm{y}^{k}\rangle.$ (A.6)
By summing (A.4), (A.5) and (A.6), we have that
$\;\|\bm{x}^{k+1}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k+1}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{*}\|^{2}$ (A.7)
$=$ $\;\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}$
$\;-2\sigma_{k}\left(\langle\bm{y}^{k+1}-\bm{y}^{*},\,\bm{\theta}^{k+1}\rangle+\langle\bm{x}^{k+1}-\bm{x}^{*},\,\bm{\xi}^{k+1}\rangle\right)-2\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle$
$\;+\|\sigma_{k}\Delta^{k+1}\|^{2}-\|\bm{x}^{k+1}-\bm{x}^{k}\|^{2}-\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{k}\|^{2}.$
Note from (A.1) that
$\big(\bm{\theta}^{k+1},\,\bm{\xi}^{k+1}\big)\in\partial\ell(\bm{y}^{k+1},\bm{x}^{k+1}),$ (A.8)
which, together with $\bm{0}\in\partial\ell(\bm{y}^{*},\bm{x}^{*})$ and the monotonicity of $\partial\ell$, yields
$\langle\bm{y}^{k+1}-\bm{y}^{*},\,\bm{\theta}^{k+1}\rangle+\langle\bm{x}^{k+1}-\bm{x}^{*},\,\bm{\xi}^{k+1}\rangle\geq 0.$
Moreover, by using (A.2), we see that
$-2\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle+\|\sigma_{k}\Delta^{k+1}\|^{2}$ $\leq 2\big|\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle\big|+\|\sigma_{k}\Delta^{k+1}\|^{2}$
$\leq\rho\left(\|\bm{x}^{k+1}-\bm{x}^{k}\|^{2}+\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{k}\|^{2}\right).$
Substituting the above two inequalities into (A.7), we obtain a key inequality for the subsequent convergence analysis:
$\;\|\bm{x}^{k+1}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k+1}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{*}\|^{2}$ (A.9)
$\leq$ $\;\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}$
$\;-(1-\rho)\left(\|\bm{x}^{k+1}-\bm{x}^{k}\|^{2}+\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{k}\|^{2}\right).$
The inequality (A.9), together with $\rho\in[0,1)$ and $\tau_{k+1}\leq(1+\nu_{k})\tau_{k}$ with $\nu_{k}\geq 0$ and $\sum\nu_{i}<\infty$ for all $k\geq 0$, implies that
$\;\|\bm{x}^{k+1}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k+1}-\bm{y}^{*}\|^{2}+\tau_{k+1}\|\bm{y}^{k+1}-\bm{y}^{*}\|^{2}$ (A.10)
$\leq$ $\;(1+\nu_{k})\left(\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right).$
Since $\{\nu_{k}\}$ is a non-negative summable sequence, it then follows from [35, Lemma 2 in Section 2.2 on Page 44] that the sequence $\left\{\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$ is convergent (but do not necessarily converge to 0). This together with $\tau_{k}\geq\tau_{\min}>0$ implies that all sequences $\{\bm{x}^{k}\}$, $\{\bm{w}^{k}\}$, $\{\bm{y}^{k}\}$ are bounded.
Statement (ii). Using (A.9) again with $\tau_{k+1}\leq(1+\nu_{k})\tau_{k}$ with $\nu_{k}\geq 0$ and $\sum\nu_{i}<\infty$ for all $k\geq 0$, we have that
$0\leq$ $\;(1-\rho){(1+\nu_{k})}\left(\|\bm{x}^{k+1}-\bm{x}^{k}\|^{2}+\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{k}\|^{2}\right)$ (A.11)
$\leq$ $\;(1+\nu_{k})\left(\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right)$
$\quad-\left(\|\bm{x}^{k+1}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k+1}-\bm{y}^{*}\|^{2}+\tau_{k+1}\|\bm{y}^{k+1}-\bm{y}^{*}\|^{2}\right).$
Since $\left\{\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$ is convergent, $\nu_{k}\to 0$ (due to $\nu_{k}\geq 0$ and $\sum\nu_{i}<\infty$) and $\rho\in[0,1)$, it then follows from (A.11) that
$\lim\limits_{k\to\infty}\|\bm{x}^{k+1}-\bm{x}^{k}\|=0\quad\mathrm{and}\quad\lim\limits_{k\to\infty}\|\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{k})\|=0.$ (A.12)
Note that both sequences $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ are bounded away from 0. Thus, we further have that $\lim\limits_{k\to\infty}\bm{\xi}^{k+1}(:=\sigma_{k}^{-1}(\bm{x}^{k}-\bm{x}^{k+1}))=\bm{0}$ and $\lim\limits_{k\to\infty}\|\bm{y}^{k+1}-\bm{y}^{k}\|=\bm{0}$. Moreover, using (A.12) together with (A.2) implies that
$\lim\limits_{k\to\infty}~|\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle|=0\quad\mathrm{and}\quad\lim\limits_{k\to\infty}~\|\sigma_{k}\Delta^{k+1}\|^{2}=0.$
Since $\{\sigma_{k}\}$ is bounded away from 0, we then obtain that $\lim\limits_{k\to\infty}\langle\bm{w}^{k}-\bm{y}^{k+1},\,\Delta^{k+1}\rangle=0$ and $\lim\limits_{k\to\infty}\Delta^{k+1}=\bm{0}$. Finally, recall again that $\tau_{k+1}\leq(1+\nu_{k})\tau_{k}$ with $\nu_{k}\geq 0$ and $\sum\nu_{i}<\infty$ for all $k\geq 0$. Thus, $\tau_{k}$ must be bounded from above and hence $\tau_{k}\sigma_{k}^{-1}$ is also bounded from above. Consequently, we can obtain that $\lim\limits_{k\to\infty}\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k+1}-\bm{y}^{k})=\bm{0}$ and hence $\lim\limits_{k\to\infty}\bm{\theta}^{k+1}=\bm{0}$.
Statement (iii). We first study the limit of $\{F\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)\}$. Using relations (A.8) and (3.3), we have that $(\bm{y}^{k+1},\bm{\xi}^{k+1})\in\partial F\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)$. Then, by the concavity of $F$, it holds that, for all $k\geq 0$,
$F(\bm{0},\,\bm{x}^{*})\leq F\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)+\langle\bm{y}^{k+1},\,\bm{\theta}^{k+1}\rangle+\langle\bm{\xi}^{k+1},\,\bm{x}^{k+1}-\bm{x}^{*}\rangle.$
Since $\lim\limits_{k\to\infty}\bm{\theta}^{k+1}=\bm{0}$, $\lim\limits_{k\to\infty}\bm{\xi}^{k+1}=\bm{0}$, and the sequences $\{\bm{x}^{k}\}$ and $\{\bm{y}^{k}\}$ are bounded, we can obtain from the above inequality that
$\liminf\limits_{k\to\infty}\,F\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)\geq F(\bm{0},\,\bm{x}^{*}).$ (A.13)
On the other hand, since $\{\bm{x}^{k}\}$ is bounded, it has at least one accumulation point. Suppose that $\bm{x}^{\infty}$ is an accumulation point and $\{\bm{x}^{k_{i}}\}$ is a convergent subsequence such that $\lim\limits_{i\to\infty}\bm{x}^{k_{i}}=\bm{x}^{\infty}$. Since $\lim\limits_{k\to\infty}\|\bm{x}^{k+1}-\bm{x}^{k}\|=0$, we also have that $\lim\limits_{i\to\infty}\bm{x}^{k_{i}+1}=\bm{x}^{\infty}$. Thus, by passing to a further subsequence if necessary, we may assume without loss of generality that the subsequence $\big\{F\big(\bm{\theta}^{k_{i}+1},\,\bm{x}^{k_{i}+1}\big)\big\}$ satisfies
$\lim\limits_{i\to\infty}F\big(\bm{\theta}^{k_{i}+1},\bm{x}^{k_{i}+1}\big)=\limsup\limits_{k\to\infty}\,F\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big).$
Note that $F$ is closed upper semicontinuous concave (see, for example, [40, Theorem 7]), and thus $\mathrm{dom}\,F$ is closed. This together with $(\bm{0},\,\bm{x}^{\infty})=\lim\limits_{i\to\infty}\big(\bm{\theta}^{k_{i}+1},\,\bm{x}^{k_{i}+1}\big)$ induces that $(\bm{0},\,\bm{x}^{\infty})\in\mathrm{dom}\,F$. Then, we see that
$\;\;\;\;\;F(\bm{0},\,\bm{x}^{*})$
$\geq F(\bm{0},\,\bm{x}^{\infty})$ (since $\bm{x}^{*}$ is optimal for problem (3.2))
$=F\bigg(\lim\limits_{i\to\infty}\bm{\theta}^{k_{i}+1},\,\lim\limits_{i\to\infty}\bm{x}^{k_{i}+1}\bigg)$ (since $\lim\limits_{k\to\infty}\bm{\theta}^{k+1}=\bm{0}$ and $\lim\limits_{i\to\infty}\bm{x}^{k_{i}+1}=\bm{x}^{\infty}$)
$\geq\limsup\limits_{i\to\infty}\,F\big(\bm{\theta}^{k_{i}+1},\,\bm{x}^{k_{i}+1}\big)$ (since $F$ is upper semicontinuous)
$=\limsup\limits_{k\to\infty}\,F\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big).$ (by the choice of subsequence $\{\bm{x}^{k_{i}+1}\}$)
This together with (A.13) implies that
$\lim\limits_{k\to\infty}\,F\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)=F(\bm{0},\,\bm{x}^{*}).$ (A.14)
We next study the limit of $\{G(\bm{y}^{k+1},\bm{\xi}^{k+1})\}$. Since F-F and $G$ are convex conjugate and $\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)\in\partial G(\bm{y}^{k+1},\,\bm{\xi}^{k+1})$, we can get the following equality by using the Fenchel equality (see, for example, [39, Theorem 23.5]):
$G(\bm{y}^{k+1},\,\bm{\xi}^{k+1})=F\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)+\langle\bm{\theta}^{k+1},\,\bm{y}^{k+1}\rangle+\langle\bm{x}^{k+1},\,\bm{\xi}^{k+1}\rangle.$
Since $\lim\limits_{k\to\infty}\bm{\theta}^{k+1}=\bm{0}$, $\lim\limits_{k\to\infty}\bm{\xi}^{k+1}=\bm{0}$, and the sequences $\{\bm{x}^{k}\}$ and $\{\bm{y}^{k}\}$ are bounded, we obtain that
$\lim\limits_{k\to\infty}\,G(\bm{y}^{k+1},\bm{\xi}^{k+1})=F(\bm{0},\,\bm{x}^{*})=G(\bm{y}^{*},\,\bm{0}).$
This proves statement (iii).
Statement (iv). We first prove that any accumulation point of $\{\bm{y}^{k}\}$ is an optimal solution of problem (2.1). Since $\{\bm{y}^{k}\}$ is bounded by statement (i), the sequence $\{\bm{y}^{k}\}$ has at least one accumulation point. Suppose that $\bm{y}^{\infty}$ is an accumulation point and $\{\bm{y}^{k_{j}}\}$ is a convergent subsequence such that $\lim\limits_{j\to\infty}\bm{y}^{k_{j}}=\bm{y}^{\infty}$. Since $\lim\limits_{k\to\infty}\|\bm{y}^{k+1}-\bm{y}^{k}\|=0$, we also have that $\lim\limits_{j\to\infty}\bm{y}^{k_{j}+1}=\bm{y}^{\infty}$. Then, using the fact that $G$ is lower semicontinuous and convex, and $\lim\limits_{k\to\infty}\bm{\xi}^{k+1}=\bm{0}$, we obtain that
$G(\bm{y}^{\infty},\bm{0})=G(\lim_{j\to\infty}\bm{y}^{k_{j}+1},\,\lim_{j\to\infty}\bm{\xi}^{k_{j}+1})\leq\liminf_{j\to\infty}~G(\bm{y}^{k_{j}+1},\,\bm{\xi}^{k_{j}+1})=G(\bm{y}^{*},\bm{0}).$
This implies that $\bm{y}^{\infty}$ is an optimal solution of problem (2.1). Similarly, using the upper semicontinuity of $F$ and analogous manipulations, we can prove that any accumulation point of $\{\bm{x}^{k}\}$ is an optimal solution of problem (3.2). This proves statement (iv).
Statement (v). We next prove that the whole sequence $\{\bm{x}^{k}\}$ is convergent. Let
$D_{\tau_{k}}\left((\bm{w}^{k},\bm{y}^{k}),\,\mathcal{Y}^{*}\right):=\inf\limits_{\bm{y}^{*}\in\mathcal{Y}^{*}}\left\{\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\},$
and define that
$\phi:=\liminf_{k\to\infty}\,D_{\tau_{k}}\left((\bm{w}^{k},\bm{y}^{k}),\,\mathcal{Y}^{*}\right),$
where $\mathcal{Y}^{*}$ is the solution set of problem (2.1) (i.e., problem (1.2)). Since $\{\bm{w}^{k}\}$, $\{\bm{y}^{k}\}$ and $\{\tau_{k}\}$ are bounded, we see that $0\leq\phi<\infty$ and there exists a subsequence $\{(\bm{w}^{k_{j}},\bm{y}^{k_{j}},\tau_{k_{j}})\}$ such that
$\lim_{j\to\infty}\,D_{\tau_{k_{j}}}\left((\bm{w}^{k_{j}},\bm{y}^{k_{j}}),\,\mathcal{Y}^{*}\right)=\phi.$
Then, by passing to a further subsequence if necessary, we may also assume without loss of generality that the subsequence $\{\bm{x}^{k_{j}}\}\subseteq\{\bm{x}^{k}\}$ converges to some accumulation point $\bm{x}^{\infty}$, which, in view of statement (iv), belongs to $\mathcal{X}^{*}$ (the optimal solution set of (3.2)). Thus, for such $\bm{x}^{\infty}$ and any $\bm{y}^{*}\in\mathcal{Y}^{*}$, using (A.10) with some manipulations, we can obtain that, for all $k>k_{j}$,
$\,\|\bm{x}^{k}-\bm{x}^{\infty}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}$
$\leq$ $\;\left(\prod_{i=k_{j}}^{k-1}\big(1+\nu_{i}\big)\right)\Big(\|\bm{x}^{k_{j}}-\bm{x}^{\infty}\|^{2}+\|\bm{w}^{k_{j}}-\bm{y}^{*}\|^{2}+\tau_{k_{j}}\|\bm{y}^{k_{j}}-\bm{y}^{*}\|^{2}\Big).$
Since $0\leq\phi\leq\liminf\limits_{k\to\infty}\left\{\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$ for any $\bm{y}^{*}\in\mathcal{Y}^{*}$, passing to the limit superior when $k\to\infty$ on the both sides of the above inequality, we obtain that, for any $\bm{y}^{*}\in\mathcal{Y}^{*}$,
$\;\phi+\limsup\limits_{k\to\infty}\left\{\|\bm{x}^{k}-\bm{x}^{\infty}\|^{2}\right\}$
$\leq$ $\;\limsup\limits_{k\to\infty}\left\{\|\bm{x}^{k}-\bm{x}^{\infty}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$
$\leq$ $\;\left(\prod_{i=k_{j}}^{\infty}(1+\nu_{i})\right)\left(\|\bm{x}^{k_{j}}-\bm{x}^{\infty}\|^{2}+\|\bm{w}^{k_{j}}-\bm{y}^{*}\|^{2}+\tau_{k_{j}}\|\bm{y}^{k_{j}}-\bm{y}^{*}\|^{2}\right),\quad\forall j\geq 0.$
Taking the infimum in $\bm{y}^{*}\in\mathcal{Y}^{*}$ on the right-hand side of the last inequality, we have that
$\limsup\limits_{k\to\infty}\left\{\|\bm{x}^{k}-\bm{x}^{\infty}\|^{2}\right\}$
$\leq$ $\;\left(\prod_{i=k_{j}}^{\infty}(1+\nu_{i})\right)\|\bm{x}^{k_{j}}-\bm{x}^{\infty}\|^{2}+\left(\prod_{i=k_{j}}^{\infty}(1+\nu_{i})\right)D_{\tau_{k_{j}}}\left((\bm{w}^{k_{j}},\bm{y}^{k_{j}}),\,\mathcal{Y}\right)-\phi,\quad\forall j\geq 0.$
Since $\ln\left(\prod_{i=k_{j}}^{\infty}(1+\nu_{i})\right)=\sum_{i=k_{j}}^{\infty}\ln(1+\nu_{i})\leq\sum_{i=k_{j}}^{\infty}\nu_{i}$ and $\lim\limits_{j\to\infty}\sum_{i=k_{j}}^{\infty}\nu_{i}=0$ (due to the summability of $\{\nu_{k}\}$), we see that $\lim\limits_{j\to\infty}\prod_{i=k_{j}}^{\infty}(1+\nu_{i})=1$. Using this fact, we can observe that the right-hand side of the above inequality converges to 0 as $j\to\infty$. Then, we conclude that $\lim\limits_{k\to\infty}\bm{x}^{k}=\bm{x}^{\infty}$, which completes the proof.    $\Box$
A.2 Proof of Theorem 3.2
Proof.
Statement (i). For the sake of clarity, we will present our proof in three steps.
Step I. Let $(\bm{y}^{*},\bm{x}^{*})\in\mathbb{R}^{M}\times\mathbb{R}^{N}$ be an arbitrary saddle point of $\ell$. Similar to the proof of Theorem 3.1(i), we combine (A.4) with (A.6) to obtain that
$\;\|\bm{x}^{k+1}-\bm{x}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{*}\|^{2}$
$=$ $\;\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\;-2\sigma_{k}\left(\langle\bm{y}^{k+1}-\bm{y}^{*},\,\bm{\theta}^{k+1}\rangle+\langle\bm{x}^{k+1}-\bm{x}^{*},\,\bm{\xi}^{k+1}\rangle\right)$
$\;+2\sigma_{k}\langle\bm{y}^{k+1}-\bm{y}^{*},\Delta^{k+1}\rangle-\|\bm{x}^{k+1}-\bm{x}^{k}\|^{2}-\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{k}\|^{2}.$
Since $(\bm{0},\bm{0})\in\partial\ell(\bm{y}^{*},\bm{x}^{*})$ and $\left(\bm{\theta}^{k+1},\,\bm{\xi}^{k+1}\right)\in\partial\ell(\bm{y}^{k+1},\bm{x}^{k+1})$, it then follows from the monotonicity of $\partial\ell$ that
$\langle\bm{y}^{k+1}-\bm{y}^{*},\,\bm{\theta}^{k+1}\rangle+\langle\bm{x}^{k+1}-\bm{x}^{*},\,\bm{\xi}^{k+1}\rangle\geq 0.$
Thus, we conclude that
$\;\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k}-\bm{y}^{*})\\ \bm{x}^{k}-\bm{x}^{*}\end{matrix}\right\|^{2}-\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{*})\\ \bm{x}^{k+1}-\bm{x}^{*}\end{matrix}\right\|^{2}$ (A.15)
$\geq$ $\;\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{k})\\ \bm{x}^{k+1}-\bm{x}^{k}\end{matrix}\right\|^{2}-2\sigma_{k}\|\bm{y}^{k+1}-\bm{y}^{*}\|\|\Delta^{k+1}\|.$
Define the sequences $\{\overline{\bm{y}}^{k}\}\subseteq\mathbb{R}^{M}$ and $\{\overline{\bm{x}}^{k}\}\subseteq\mathbb{R}^{N}$ as follows:
$\overline{\bm{y}}^{k}:=\Pi_{\mathcal{Y}^{*}}(\bm{y}^{k})\quad\text{and}\quad\overline{\bm{x}}^{k}:=\Pi_{\mathcal{X}^{*}}(\bm{x}^{k}),\quad\forall\;k\geq 0,$
where $\mathcal{Y}^{*}$ is the solution set of problem (2.1) (i.e., problem (1.2)), $\mathcal{X}^{*}$ is the solution set of problem (3.2) (i.e., problem (1.1)), and $\Pi_{\mathcal{Y}^{*}}(\bm{y}^{k})$ (resp. $\Pi_{\mathcal{X}^{*}}(\bm{x}^{k})$) denotes the projection of $\bm{y}^{k}$ (resp. $\bm{x}^{k}$) onto set $\mathcal{Y}^{*}$ (resp. $\mathcal{X}^{*}$). Since (A.15) holds for any $\bm{y}^{*}\in\mathcal{Y}^{*}$ and $\bm{x}^{*}\in\mathcal{X}^{*}$, we can replace $\bm{y}^{*}$ and $\bm{x}^{*}$ with $\overline{\bm{y}}^{k}$ and $\overline{\bm{x}}^{k}$, respectively, to obtain that
$\;\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k}-\overline{\bm{y}}^{k})\\ \bm{x}^{k}-\overline{\bm{x}}^{k}\end{matrix}\right\|^{2}-\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\overline{\bm{y}}^{k})\\ \bm{x}^{k+1}-\overline{\bm{x}}^{k}\end{matrix}\right\|^{2}$
$\geq$ $\;\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{k})\\ \bm{x}^{k+1}-\bm{x}^{k}\end{matrix}\right\|^{2}-2\sigma_{k}\|\bm{y}^{k+1}-\overline{\bm{y}}^{k}\|\|\Delta^{k+1}\|.$
Moreover, from the definitions of $\overline{\bm{y}}^{k}$ and $\overline{\bm{x}}^{k}$, we have that $\left\|\bm{y}^{k+1}-\overline{\bm{y}}^{k+1}\right\|\leq\left\|\bm{y}^{k+1}-\overline{\bm{y}}^{k}\right\|$ and $\left\|\bm{x}^{k+1}-\overline{\bm{x}}^{k+1}\right\|\leq\left\|\bm{x}^{k+1}-\overline{\bm{x}}^{k}\right\|$. These, together with the above inequality, yield that
$\;\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k}-\overline{\bm{y}}^{k})\\ \bm{x}^{k}-\overline{\bm{x}}^{k}\end{matrix}\right\|^{2}-\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\overline{\bm{y}}^{k+1})\\ \bm{x}^{k+1}-\overline{\bm{x}}^{k+1}\end{matrix}\right\|^{2}$ (A.16)
$\geq$ $\;\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{k})\\ \bm{x}^{k+1}-\bm{x}^{k}\end{matrix}\right\|^{2}-2\sigma_{k}\|\bm{y}^{k+1}-\overline{\bm{y}}^{k}\|\|\Delta^{k+1}\|.$
Step II. We next derive an upper bound for $\|\bm{y}^{k+1}-\overline{\bm{y}}^{k}\|\|\Delta^{k+1}\|$. On the one hand, we have from (A.2) that
$\sigma_{k}^{2}\|\Delta^{k+1}\|^{2}\leq\rho\left(\|\bm{x}^{k+1}-\bm{x}^{k}\|^{2}+\tau_{k}\|\bm{y}^{k+1}-\bm{y}^{k}\|^{2}\right),$
which implies that
$\|\Delta^{k+1}\|\leq\frac{\sqrt{\rho}}{\sigma_{k}}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{k})\\ \bm{x}^{k+1}-\bm{x}^{k}\end{matrix}\right\|.$ (A.17)
On the other hand, we see that
$\|\bm{y}^{k+1}-\overline{\bm{y}}^{k}\|$ $\leq\|\bm{y}^{k+1}-\overline{\bm{y}}^{k+1}\|+\|\overline{\bm{y}}^{k+1}-\overline{\bm{y}}^{k}\|\leq\|\bm{y}^{k+1}-\overline{\bm{y}}^{k+1}\|+\|\bm{y}^{k+1}-{\bm{y}}^{k}\|$ (A.18)
$\leq\|\bm{y}^{k+1}-\overline{\bm{y}}^{k+1}\|+\frac{1}{\sqrt{\tau_{k}}}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-{\bm{y}}^{k})\\ \bm{x}^{k+1}-{\bm{x}}^{k}\end{matrix}\right\|,$
where the second inequality follows from the non-expansiveness of the projection operator $\Pi_{\mathcal{Y}^{*}}(\cdot)$. Moreover, since $\{\bm{y}^{k}\}$ and $\{\bm{x}^{k}\}$ are bounded (by Theorem 3.1(i)), there must exist a positive scalar $r$ such that
$\mathrm{dist}\big((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\big)\leq r,\quad\forall\,k\geq 0.$
Thus, we apply Assumption A with this $r$ and know that, there exists a $\kappa>0$ such that
$\operatorname{dist}\left((\bm{y}^{k+1},\bm{x}^{k+1}),\,\mathcal{Y}^{*}\times\mathcal{X}^{*}\right)\leq\kappa\operatorname{dist}\left((\bm{0},\bm{0}),\,\partial\ell(\bm{y}^{k+1},\bm{x}^{k+1})\right)\leq\kappa\left\|(\bm{\theta}^{k+1},\,\bm{\xi}^{k+1})\right\|,$
where the last inequality is due to (A.8) (i.e., $\left(\bm{\theta}^{k+1},\,\bm{\xi}^{k+1}\right)\in\partial\ell(\bm{y}^{k+1},\bm{x}^{k+1})$) with $\bm{\theta}^{k+1}:=\Delta^{k+1}-\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k+1}-\bm{y}^{k})$ and $\bm{\xi}^{k+1}:=\sigma_{k}^{-1}(\bm{x}^{k}-\bm{x}^{k+1})$. This inequality further implies that
$\|\bm{y}^{k+1}-\overline{\bm{y}}^{k+1}\|\leq\;\sqrt{\|\bm{y}^{k+1}-\overline{\bm{y}}^{k+1}\|^{2}+\|\bm{x}^{k+1}-\overline{\bm{x}}^{k+1}\|^{2}}$ (A.19)
$\leq$ $\;\kappa\left\|\begin{matrix}\Delta^{k+1}-\tau_{k}\sigma_{k}^{-1}(\bm{y}^{k+1}-\bm{y}^{k})\\ \sigma_{k}^{-1}(\bm{x}^{k}-\bm{x}^{k+1})\end{matrix}\right\|\;\leq\;\kappa\left(\|\Delta^{k+1}\|+\frac{1}{\sigma_{k}}\left\|\begin{matrix}\tau_{k}(\bm{y}^{k+1}-\bm{y}^{k})\\ \bm{x}^{k+1}-\bm{x}^{k}\end{matrix}\right\|\right)$
$\leq$ $\;\kappa\left(\|\Delta^{k+1}\|+\frac{\sqrt{\overline{\tau}_{k}}}{\sigma_{k}}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{k})\\ \bm{x}^{k+1}-\bm{x}^{k}\end{matrix}\right\|\right)$
$\leq$ $\;\frac{\kappa\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{k}}\right)}{\sigma_{k}}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{k})\\ \bm{x}^{k+1}-\bm{x}^{k}\end{matrix}\right\|,$
where $\overline{\tau}_{k}$ is defined as $\overline{\tau}_{k}:=\max\left\{1,\tau_{k}\right\}$ and the last inequality follows from (A.17). Now, combining (A.17), (A.18) and (A.19), with some manipulations, we can obtain that
$\|\bm{y}^{k+1}-\overline{\bm{y}}^{k}\|\|\Delta^{k+1}\|\leq\left(\frac{\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)+\sigma_{k}\sqrt{\rho}}{\sigma_{k}^{2}\sqrt{\tau_{k}}}\right)\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{k})\\ \bm{x}^{k+1}-\bm{x}^{k}\end{matrix}\right\|^{2}.$
Then, substituting this inequality into (A.16) yields that
$\;\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k}-\overline{\bm{y}}^{k})\\ \bm{x}^{k}-\overline{\bm{x}}^{k}\end{matrix}\right\|^{2}-\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\overline{\bm{y}}^{k+1})\\ \bm{x}^{k+1}-\overline{\bm{x}}^{k+1}\end{matrix}\right\|^{2}$ (A.20)
$\geq$ $\;\left(1-\frac{2\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)+2\sigma_{k}\sqrt{\rho}}{\sigma_{k}\sqrt{\tau_{k}}}\right)\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-{\bm{y}}^{k})\\ \bm{x}^{k+1}-{\bm{x}}^{k}\end{matrix}\right\|^{2}.$
Step III. In the following, we will establish the convergence rate based on (A.20). First, by recalling the conditions on $\{\tau_{k}\}$: $\tau_{k}\geq\tau_{\min}>0$ and $\tau_{k+1}\leq(1+\nu_{k})\tau_{k}$ with $\nu_{k}\geq 0$ and $\sum_{k=0}^{\infty}\nu_{k}<+\infty$, we know that there exists $\tau_{\max}:=\left(\prod_{k=0}^{\infty}(1+\nu_{k})\right)\tau_{0}$ such that $0<\tau_{\min}\leq\tau_{k}\leq\tau_{\max}<+\infty$ for all $k\geq 0$. This together with condition (3.5) implies that there exists a positive integer $k_{0}$ such that
$\sqrt{\tau_{k}}-2\sqrt{\rho}>0\quad\text{and}\quad\sigma_{k}>c\cdot\frac{2\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)}{\sqrt{\tau_{k}}-2\sqrt{\rho}},\quad\forall\,k\geq k_{0},$
where $c>1$. Hence, one can verify that
$\left(1-\frac{2\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)+2\sigma_{k}\sqrt{\rho}}{\sigma_{k}\sqrt{\tau_{k}}}\right)>\widetilde{c}:=\Big(\frac{c-1}{c}\Big)\cdot\frac{\sqrt{\tau_{\min}}-2\sqrt{\rho}}{\sqrt{\tau_{\min}}}>0,\quad\forall\,k\geq k_{0},$ (A.21)
which means that the factor in the right-hand side of (A.20) will be positive when $k\geq k_{0}$. On the other hand, using (A.19) again, we deduce that
$\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\bm{y}^{k})\\ \bm{x}^{k+1}-\bm{x}^{k}\end{matrix}\right\|^{2}$ $\geq\frac{\sigma_{k}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{k}}\right)^{2}}\left\|\begin{matrix}\bm{y}^{k+1}-\overline{\bm{y}}^{k+1}\\ \bm{x}^{k+1}-\overline{\bm{x}}^{k+1}\end{matrix}\right\|^{2}$
$\geq\frac{\sigma_{k}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{k}}\right)^{2}\overline{\tau}_{k}}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\overline{\bm{y}}^{k+1})\\ \bm{x}^{k+1}-\overline{\bm{x}}^{k+1}\end{matrix}\right\|^{2}.$
This, together with (A.20) and (A.21), yields that
$\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k}-\overline{\bm{y}}^{k})\\ \bm{x}^{k}-\overline{\bm{x}}^{k}\end{matrix}\right\|^{2}\geq\left(1+\gamma_{k}\right)\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-\overline{\bm{y}}^{k+1})\\ \bm{x}^{k+1}-\overline{\bm{x}}^{k+1}\end{matrix}\right\|^{2},\quad\forall\,k\geq k_{0},$ (A.22)
where
$\gamma_{k}\,$= $\,\left(1-\frac{2\kappa\sqrt{\tau_{k}}\left(\rho+\sqrt{\rho\overline{\tau}_{k}}\right)+2\sigma_{k}\sqrt{\rho}}{\sigma_{k}\sqrt{\tau_{k}}}\right)\frac{\sigma_{k}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{k}}\right)^{2}\overline{\tau}_{k}}$ (A.23)
$\geq$ $\,\widetilde{c}\cdot\frac{\sigma_{k}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{\max}}\right)^{2}\overline{\tau}_{\max}}$
$\geq$ $\,\gamma_{\min}=\frac{\widetilde{c}\,\sigma_{\min}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{\max}}\right)^{2}\overline{\tau}_{\max}}>0,\quad\forall\,k\geq k_{0}.$
Let $\Lambda^{k}:=\operatorname{Diag}(\tau_{k}I_{M},I_{N})$. Since $\{\tau_{k}\}$ is bounded away from 0 and satisfies that $(1+\nu_{k})\tau_{k}\geq\tau_{k+1}$, we have that $(1+\nu_{k})\Lambda^{k}\succeq\Lambda^{k+1}\succ 0$. Then, one can obtain from (A.22) that
$(1+\nu_{k})\left\|\begin{matrix}\bm{y}^{k}-\overline{\bm{y}}^{k}\\ \bm{x}^{k}-\overline{\bm{x}}^{k}\end{matrix}\right\|_{\Lambda^{k}}^{2}\geq\left(1+\gamma_{k}\right)\left\|\begin{matrix}\bm{y}^{k+1}-\overline{\bm{y}}^{k+1}\\ \bm{x}^{k+1}-\overline{\bm{x}}^{k+1}\end{matrix}\right\|_{\Lambda^{k+1}}^{2},$
which readily implies that
$\mathrm{dist}_{\Lambda^{k+1}}\left((\bm{y}^{k+1},\bm{x}^{k+1}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right)\leq\mu_{k}\,\mathrm{dist}_{\Lambda^{k}}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right),$
where $\mu_{k}:=\sqrt{\frac{1+\nu_{k}}{1+\gamma_{k}}}$. Since $\nu_{k}\to 0$ and $\gamma_{k}\geq\gamma_{\min}>0$ for all $k\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\geq k_{0}$, one can verify that $\limsup\limits_{k\to\infty}\,\{\mu_{k}\}<1$. Thus, we obtain the desired results in statement (i).
Statement (ii). Using (A.20) and (A.21) again, we see that
$\widetilde{c}\,\tau_{\min}\|\bm{y}^{k+1}-{\bm{y}}^{k}\|^{2}\leq\widetilde{c}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-{\bm{y}}^{k})\\ \bm{x}^{k+1}-{\bm{x}}^{k}\end{matrix}\right\|^{2}\leq\mathrm{dist}_{\Lambda^{k}}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}2}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right),$
for any $k\geq k_{0}$. Using this inequality and the fact that the sequence $\left\{\mathrm{dist}_{\Lambda^{k}}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right)\right\}$ is asymptotically Q-(super)linear convergent, we can conclude that there exist a positive integer $k_{1}$, $0<\beta<1$ and $C>0$ such that
$\|\bm{y}^{k+1}-{\bm{y}}^{k}\|\leq C\beta^{k},\quad\forall\,k\geq k_{1},$
which further implies that $\sum_{k=0}^{\infty}\|\bm{y}^{k+1}-{\bm{y}}^{k}\|<\infty$. Consequently, $\{\bm{y}^{k}\}$ is a Cauchy sequence and hence convergent. Therefore, the proof is completed.    $\Box$
Appendix B A dual ADMM for QROT
In this section, we present the alternating direction method of multipliers (ADMM, see, e.g. [6, 17]) for solving the dual problem (4.3), which can be reformulated as:
$\min\limits_{\bm{u}\in\mathbb{R}^{m},\,\bm{v}\in\mathbb{R}^{n},\,W\in\mathbb{R}^{m\times n}}\Big\{f_{\texttt{q}}^{*}(W)-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v}\;\mid\;\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}=W\Big\}.$ (B.1)
Given a penalty parameter $\sigma>0$, the augmented Lagrangian function of (B.1) is
$\;\mathcal{L}_{\sigma}\left(\bm{u},\bm{v},W;X\right)$
= $\;f_{\texttt{q}}^{*}(W)-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v}+\big\langle X,\,\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-W\big\rangle+\frac{\sigma}{2}\big\|\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-W\big\|_{F}^{2}.$
Then, the ADMM for solving (B.1) can be described as in Algorithm 3.
Algorithm 3 ADMM for solving (B.1)
Input: a penalty parameter $\sigma>0$, and initializations $\bm{u}^{0}\in\mathbb{R}^{m}$, $\bm{v}^{0}\in\mathbb{R}^{n}$, $W^{0},X^{0}\in\mathbb{R}^{m\times n}$. Set $k=0$.
while the termination criterion is not met, do
  Step 1. Compute
$\big(\bm{u}^{k+1},\,\bm{v}^{k+1}\big)=\arg\min\limits_{\bm{u},\bm{v}}~\mathcal{L}_{\sigma}\big(\bm{u},\bm{v},W^{k};X^{k}\big).$
  Step 2. Compute
$W^{k+1}=\arg\min\limits_{W}~\mathcal{L}_{\sigma}\big(\bm{u}^{k+1},\bm{v}^{k+1},W;X^{k}\big).$
  Step 3. Set $X^{k+1}=X^{k}+\tau\sigma\left(\bm{u}^{k+1}\bm{1}_{n}^{\top}+\bm{1}_{m}(\bm{v}^{k+1})^{\top}-W^{k+1}\right)$, where $\tau\in\left(0,\frac{1+\sqrt{5}}{2}\right)$ is the dual step-size that is typically set to $1.618$.
  Step 4. Set $k\leftarrow k+1$ and go to Step 1.
end while
Output: $\big(\bm{u}^{k},\bm{v}^{k},W^{k},X^{k}\big)$.
Both subproblems in ADMM can be solved efficiently. Specifically, $\big(\bm{u}^{k+1},\,\bm{v}^{k+1}\big)$ can be obtained by solving the following unconstrained convex minimization problem:
$\min\limits_{\bm{u},\bm{v}}~h_{k}(\bm{u},\bm{v}):=-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v}+\frac{\sigma}{2}\left\|\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-S^{k}\right\|_{F}^{2},$ (B.2)
where $S^{k}:=W^{k}-\sigma^{-1}X^{k}$. From the first-order optimality conditions of (B.2), we see that solving problem (B.2) is equivalent to solving the equation $\nabla h_{k}(\bm{u},\bm{v})=\bm{0}$. This, in turn, reduces to solving the following linear system
$n\bm{u}+(\bm{1}_{n}^{\top}\bm{v})\bm{1}_{m}=\sigma^{-1}\bm{\alpha}+S^{k}\bm{1}_{n},$ (B.3)
$(\bm{1}_{m}^{\top}\bm{u})\bm{1}_{n}+m\bm{v}=\sigma^{-1}\bm{\beta}+(S^{k})^{\top}\bm{1}_{m}.$ (B.4)
With some algebraic manipulations, it is not difficult to show that
$\bm{u}^{*}(t)$ $=$ $\frac{\sigma^{-1}\bm{\alpha}+S^{k}\bm{1}_{n}}{n}+t\,\bm{1}_{m},\quad\forall\,t\in\mathbb{R},$
$\bm{v}^{*}(t)$ $=$ $\frac{\sigma^{-1}\bm{\beta}+(S^{k})^{\top}\bm{1}_{m}}{m}-\frac{\bm{1}_{m}^{\top}\bm{u}^{*}(t)}{m}\bm{1}_{n},\quad\forall\,t\in\mathbb{R}.$
solves the above linear system. Thus, we obtain a solution pair $(\bm{u}^{*}(t),\bm{v}^{*}(t))$ with any $t\in\mathbb{R}$. It can be routinely shown that $(\bm{u}^{*}(t),\bm{v}^{*}(t))$ satisfies the linear system (B.3) and (B.4), and therefore solves problem (B.2). On the other hand, $W^{k+1}$ can be obtained by computing the proximal operator of the function $\sigma^{-1}f_{\texttt{q}}^{*}$, i.e.,
$W^{k+1}~:=~\mathtt{prox}_{\sigma^{-1}f_{\texttt{q}}^{*}}\big(Q^{k}\big)=~\left\{\begin{aligned} &C-\Pi_{\mathbb{R}^{m\times n}_{+}}\big(C-Q^{k}\big),&&\lambda=0,\\[3.0pt] &Q^{k}-(1+\lambda\sigma)^{-1}\Pi_{\mathbb{R}^{m\times n}_{+}}(Q^{k}-C),&&\lambda>0,\end{aligned}\right$.
where $Q^{k}:=\bm{u}^{k+1}\bm{1}_{n}^{\top}+\bm{1}_{m}({\bm{v}}^{k+1})^{\top}+\sigma^{-1}X^{k}$.
Appendix C An iBPGM for QROT
In this section, we briefly discuss how to employ an inexact Bregman proximal gradient method (iBPGM) with Sinkhorn’s algorithm as a subsolver for solving the QROT problem (4.1). We refer readers to [59, Section 5] for more details. Specifically, the iBPGM with the entropy kernel function for solving (4.1) can be given as follows: let $X^{0}>0$ and $\phi(X):=\sum_{ij}x_{ij}(\log x_{ij}-1)$, at the $k$-th iteration, compute
$X^{k+1}$$\approx$ $\min\limits_{X}\Big\{\langle C+\lambda X^{k},\,X\rangle+\mu_{k}\mathcal{D}_{\phi}(X,\,X^{k})\;\mid\;X\bm{1}_{n}=\bm{\alpha},~~X^{\top}\bm{1}_{m}=\bm{\beta}\Big\},$ (C.1)
where $\mu_{k}\geq\lambda$ is a positive proximal parameter, and $\mathcal{D}_{\phi}(U,\,V)$ denotes the Bregman distance between $U$ and $V$ associated with the kernel function $\phi$ which is defined as $\mathcal{D}_{\phi}(U,\,V):=\phi(U)-\phi(V)-\langle\nabla\phi(V),\,U-V\rangle$. Problem (C.1) can be rewritten as
$\min\limits_{X}\Big\{\langle M^{k},\,X\rangle+\mu_{k}\,{\textstyle\sum_{ij}}x_{ij}(\log x_{ij}-1)\;\mid\;X\bm{1}_{n}=\bm{\alpha},~~X^{\top}\bm{1}_{m}=\bm{\beta}\Big\},$ (C.2)
where $M^{k}:=C+\lambda X^{k}-\mu_{k}\log X^{k}$. Note that problem (C.2) has the same form as the entropic regularized optimal transport problem and hence can be readily solved by the popular Sinkhorn’s algorithm; see [34, Section 4.2] for more details. Specifically, let $\Xi^{k}:=e^{-M^{k}/\mu_{k}}$. Then, given an initial positive vector $\bm{v}^{k,0}$, the iterative scheme is given by
$\bm{u}^{k,t}=\bm{\alpha}./\big(\Xi^{k}\bm{v}^{k,t-1}\big),\quad\bm{v}^{k,t}=\bm{\beta}./\big((\Xi^{k})^{\top}\bm{u}^{k,t}\big),\quad\forall\;t\geq 0,$ (C.3)
where ‘././’ denotes the entrywise division between two vectors. When a pair $(\bm{u}^{k,t},\,\bm{v}^{k,t})$ is obtained based on a certain stopping criterion, an approximate solution of (C.2) (and hence (C.1)) can be recovered by setting $X^{k,t}:=\mathrm{Diag}(\bm{u}^{k,t})\,\Xi^{k}\,\mathrm{Diag}(\bm{v}^{k,t})$. Meanwhile, a pair of approximate dual solutions can be recovered by setting $\bm{f}^{k,t}:=\mu_{k}\log\bm{u}^{k,t}$ and $\bm{g}^{k,t}:=\mu_{k}\log\bm{v}^{k,t}$. In our experiments, we simply execute Sinkhorn’s iteration (C.3) only once for each subproblem, and observe that this is sufficient for obtaining a promising initial point for warm-starting our ripALM.
Appendix D Efficient implementation of ripALM for BPDN
When applying ripALM to the BPDN problem (4.9), a key component is the efficient implementation of the semismooth Newton (Ssn) method for solving the non-smooth equation (4.16). The main computational task in Ssn lies in computing the Newton direction, which requires solving the linear system (4.18). In the following, we describe how this system can be solved efficiently.
Let $(\bar{\bm{s}},\bar{\bm{t}},\bm{y})\in\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{m}$ and $\sigma>0$ be given. We restate (4.18) as follows
$H\bm{d}=-\bm{g},$ (D.1)
where $\bm{g}:=\nabla\Phi(\bm{y})$, and $H\in\widehat{\partial}(\nabla\Phi)(\bm{y})$ has the form
$H:=\sigma\left(DUD^{\top}+V\right)+\frac{\tau}{\sigma}I_{m},$
with ${U}\in\partial\mathtt{prox}_{\sigma\|\cdot\|_{1}}(\bm{u})$, $\bm{u}:=\bar{\bm{s}}+\sigma{D}^{\top}\bm{y}$, and $V\in\partial\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}(\bm{v})$, $\bm{v}:=\bar{\bm{t}}-\sigma\bm{y}$. Moreover, the generalized Jacobians $U$ and $V$ are chosen as follows:
$U=$ $\;\text{Diag}(\bm{w})~~\text{with}~~w_{i}=\left\{\begin{array}[]{l}0,\quad\text{if}~|u_{i}|\leq\sigma,\\ 1,\quad\text{otherwise, }\end{array}\right.,~~i=1,\dots,n,$
[…]
Thanks to the special structure of $H$, we can solve the linear system (D.1) efficiently by considering the following two cases.
Case I. $V=I_{m}$. In this case, the linear system (D.1) reduces to
$\left(DUD^{\top}+\frac{\sigma^{2}+\tau}{\sigma^{2}}I_{m}\right)\bm{d}=-\frac{1}{\sigma}\bm{g}.$ (D.2)
Define the index set $\mathcal{I}:=\left\{i\mid|u_{i}|>\sigma,\,i=1,\ldots,n\right\}$. Taking into account the special 0-11 sparsity structure of $U$, we see that $DUD^{\top}=D_{\mathcal{I}}{D}_{\mathcal{I}}^{\top}$, where $D_{\mathcal{I}}\in\mathbb{R}^{m\times|\mathcal{I}|}$ denotes the submatrix of $D$ formed by the columns indexed by $\mathcal{I}$. Consequently, the linear system (D.2) is further simplified as
$\left(D_{\mathcal{I}}D_{\mathcal{I}}^{\top}+\bar{\gamma}I_{m}\right)\bm{d}=-\frac{1}{\sigma}\bm{g},$ (D.3)
where $\bar{\gamma}=\sigma^{-2}(\sigma^{2}+\tau)$. To solve (D.3), we proceed as follows: When $m\leq|\mathcal{I}|$, we directly factorize the coefficient matrix $D_{\mathcal{I}}D_{\mathcal{I}}^{\top}+\bar{\gamma}I_{m}$, which requires $O(m^{3})$ arithmetic operations. Otherwise, if $m>|\mathcal{I}|$, we can utilize the Sherman–Morrison–Woodbury (SMW) formula to improve efficiency. Specifically, the inverse of the coefficient matrix can be computed as
$\left(D_{\mathcal{I}}D_{\mathcal{I}}^{\top}+\bar{\gamma}I_{m}\right)^{-1}=\frac{1}{\bar{\gamma}}\left(I_{m}-D_{\mathcal{I}}\left(\bar{\gamma}I_{|\mathcal{I}|}+D_{\mathcal{I}}^{\top}D_{\mathcal{I}}\right)^{-1}D_{\mathcal{I}}^{\top}\right).$
As a result, it suffices to factorize the smaller matrix $\bar{\gamma}I_{|\mathcal{I}|}+D_{\mathcal{I}}^{\top}D_{\mathcal{I}}$, which requires $O(|\mathcal{I}|^{3})$ operations. This would be much cheaper than factorizing a matrix of size $m\times m$, when the diagonal of $U$ is sparse.
Case II. $V=\frac{\widehat{\kappa}}{\|\bm{v}\|}\left(I_{m}-\frac{1}{\|\bm{v}\|^{2}}\bm{v}\bm{v}^{\top}\right)$. In this case, the linear system (D.1) reduces to
$\left(B+D_{\mathcal{I}}D_{\mathcal{I}}^{\top}\right)\bm{d}=-\frac{1}{\sigma}\bm{g}\quad\text{with}\quad B:=\left(\frac{\widehat{\kappa}}{\|\bm{v}\|}+\frac{\tau}{\sigma^{2}}\right)I_{m}-\frac{\widehat{\kappa}}{\|\bm{v}\|^{3}}\bm{v}\bm{v}^{\top},$ (D.4)
which is more complicated than Case I due to the presence of the dense matrix $\bm{v}\bm{v}^{\top}$. When $m\leq|\mathcal{I}|$, we again directly factorize the coefficient matrix, which requires $O(m^{3})$ operations. Otherwise, if $m>|\mathcal{I}|$, we proceed as follows: We first apply the SMW formula to compute the inverse of the coefficient matrix:
$\left(B+D_{\mathcal{I}}D_{\mathcal{I}}^{\top}\right)^{-1}=B^{-1}-B^{-1}D_{\mathcal{I}}\left(I_{|\mathcal{I}|}+D_{\mathcal{I}}^{\top}B^{-1}D_{\mathcal{I}}\right)^{-1}D_{\mathcal{I}}^{\top}{B}^{-1}.$
Furthermore, the matrix ${B}^{-1}$ can be efficiently computed via the SMW formula as well:
$B^{-1}=\left(\frac{\widehat{\kappa}}{\|\bm{v}\|}+\frac{\tau}{\sigma^{2}}\right)^{-1}I_{m}+\frac{1}{\bar{\alpha}}\left(\frac{\widehat{\kappa}}{\|\bm{v}\|}+\frac{\tau}{\sigma^{2}}\right)^{-2}\frac{\widehat{\kappa}}{\|\bm{v}\|^{3}}\bm{v}\bm{v}^{\top},$
where $\bar{\alpha}:=1-\frac{\widehat{\kappa}}{\|\bm{v}\|}\left(\frac{\widehat{\kappa}}{\left\|\bm{v}\right\|}+\frac{\tau}{\sigma^{2}}\right)^{-1}>0$. Therefore, in this case, it suffices to factorize the matrix $I_{|\mathcal{I}|}+D_{\mathcal{I}}^{\top}{B}^{-1}D_{\mathcal{I}}$, which requires $O(|\mathcal{I}|^{3})$ operations.
Finally, we mention that when direct factorization of the above matrices is computationally prohibitive, one may instead apply an iterative method—such as the preconditioned conjugate gradient method—to solve the linear system (D.1). Such methods can effectively exploit the underlying sparsity and low-rank structures to reduce computational cost.
Appendix E A dual ADMM for BPDN
In this section, we briefly discuss how to apply ADMM for solving the dual problem (4.11) of the BPDN problem. To this end, we first reformulate the dual problem as:
$\min_{\bm{u},\,\bm{v},\,\bm{y}}$ $\quad\delta_{\mathcal{B}_{1}^{\infty}}(\bm{u})+\widehat{\kappa}\|\bm{v}\|-\bm{b}^{\top}\bm{y}$ (E.1)
$\mathrm{s.t.}$ $\quad D^{\top}\bm{y}-\bm{u}=0,~-\bm{y}-\bm{v}=0.$
Given a penalty parameter $\sigma>0$, the augmented Lagrangian function of (E.1) is
$\mathcal{L}_{\sigma}(\bm{u},\bm{v},\bm{y};\bm{s},\bm{t})$= $~\delta_{\mathcal{B}_{1}^{\infty}}(\bm{u})+\widehat{\kappa}\|\bm{v}\|-\bm{b}^{\top}\bm{y}+\langle\bm{s},\,D^{\top}\bm{y}-\bm{u}\rangle+\langle\bm{t},\,-\bm{y}-\bm{v}\rangle$
$+{\textstyle\frac{\sigma}{2}\left\|D^{\top}\bm{y}-\bm{u}\right\|^{2}+\frac{\sigma}{2}\left\|-\bm{y}-\bm{v}\right\|^{2}}.$
The ADMM for solving (E.1) is presented in Algorithm 4. In Step 1, we can obtain the closed-form solution to the subproblem by evaluating the proximal mappings of $\delta_{\mathcal{B}_{1}^{\infty}}(\cdot)$ and $\frac{\widehat{\kappa}}{\sigma}\|\cdot\|$. In Step 2, the subproblem reduces to solving a linear system, where the coefficient matrix $I+DD^{\top}$ is symmetric positive definite and remains unchanged throughout the iterations, and therefore requires factorization only once. For large-scale problems where such a factorization is prohibited, one may also apply the preconditioned conjugate gradient method for solving the linear system.
Algorithm 4 ADMM for solving (E.1)
Input: a penalty parameter $\sigma>0$, and initializations $\bm{y}^{0}\in\mathbb{R}^{m}$, $\bm{s}^{0}\in\mathbb{R}^{n}$, $\bm{t}^{0}\in\mathbb{R}^{m}$.
while the termination criterion is not met do
  Step 1. Compute
$\;\big(\bm{u}^{k+1},\,\bm{v}^{k+1}\big)=\arg\min_{\bm{u},\bm{v}}\left\{\mathcal{L}_{\sigma}(\bm{u},\bm{v},\bm{y}^{k};\bm{s}^{k},\bm{t}^{k})\right\}$
$\Leftrightarrow$ $\;\bm{u}^{k+1}=\Pi_{\mathcal{B}_{1}^{\infty}}\left(D^{\top}\bm{y}^{k}+\sigma^{-1}\bm{s}^{k}\right),\quad\bm{v}^{k+1}=\sigma^{-1}\left(\left(\bm{t}^{k}-\sigma\bm{y}^{k}\right)-\Pi_{{\mathcal{B}_{\widehat{\kappa}}^{2}}}\left(\bm{t}^{k}-\sigma\bm{y}^{k}\right)\right).$
  Step 2. Compute
$\;\bm{y}^{k+1}=\arg\min_{\bm{y}\in\mathbb{R}^{m}}\left\{\mathcal{L}_{\sigma}(\bm{u}^{k+1},\bm{v}^{k+1},\bm{y};\bm{s}^{k},\bm{t}^{k})\right\}$
$\Leftrightarrow$ $\;\big(DD^{\top}+I\big)\bm{y}^{k+1}=\sigma^{-1}\left(\bm{b}-D\big(\bm{s}^{k}-\sigma\bm{u}^{k+1}\big)+\big(\bm{t}^{k}-\sigma\bm{v}^{k+1}\big)\right).$
  Step 3. Compute
$\bm{s}^{k+1}=\bm{s}^{k}+\tau\sigma\big({D}^{\top}\bm{y}^{k+1}-\bm{u}^{k+1}\big),\quad\bm{t}^{k+1}=\bm{t}^{k}+\tau\sigma\big(-\bm{y}^{k+1}-\bm{v}^{k+1}\big),$
where $\tau\in\left(0,\,\frac{1+\sqrt{5}}{2}\right)$ is the dual step-size that is typically set to $1.618$.
end while
Output: $\left(\bm{s}^{k+1},\,\bm{t}^{k+1},\,\bm{y}^{k+1}\right)$.
Appendix F Second-order cone programming reformulation for BPDN
In this section, we present a second-order cone programming (SOCP) reformulation of the BPDN problem (4.9). Specifically, we introduce an auxiliary nonnegative variable $\bm{r}\in\mathbb{R}^{n}_{+}$ to majorize $|\bm{s}|$, i.e., $|\bm{s}|\leq\bm{r}$. Thus, the objective $\|\bm{s}\|_{1}$ can be replaced by $\bm{1}^{\top}\bm{r}$ together with the linear constraints $\bm{r}\leq\bm{s}\leq\bm{r}$ and $\bm{r}\geq\bm{0}$. Next, for a given positive integer $d$, we denote the $(d+1)$-dimensional second-order cone as $\mathcal{Q}^{d+1}:=\big\{(s_{0},\bm{s})\in\mathbb{R}^{d+1}\mid s_{0}\geq\|\bm{s}\|\big\}$. With these notations, the BPDN problem (4.9) admits the following SOCP reformulation:
$\min_{\bm{s},\,\bm{r}\in\mathbb{R}^{n}}$ $\quad\bm{1}^{\top}\bm{r}$
$\mathrm{s.t.}$ $\quad-\bm{r}\leq\bm{s}\leq\bm{r},\quad\bm{r}\geq\bm{0},$
$\quad\big(\widehat{\kappa},D\bm{s}-\bm{b}\big)\in\mathcal{Q}^{m+1}.$